### v1

# Harmonization in Magnetic Resonance Imaging: A Survey of Acquisition, Image-level, and Feature-level Methods

Abstract
Modern medical imaging technologies have greatly advanced neuroscience research and clinical diagnostics. However, imaging data collected across different scanners, acquisition protocols, or imaging sites often exhibit substantial heterogeneity, known as “batch effects” or “site effects.” These non-biological sources of variability can obscure true biological signals, reduce reproducibility and statistical power, and severely impair the generalizability of learning-based models across datasets. Image harmonization aims to eliminate or mitigate such site-related biases while preserving meaningful biological information, thereby improving data comparability and consistency. This review provides a comprehensive overview of key concepts, methodological advances, publicly available datasets, current challenges, and future directions in the field of medical image harmonization, with a focus on magnetic resonance imaging (MRI). We systematically cover the full imaging pipeline, and categorize harmonization approaches into prospective acquisition and reconstruction strategies, retrospective image-level and feature-level methods, and traveling-subject-based techniques. Rather than providing an exhaustive survey, we focus on representative methods, with particular emphasis on deep learning-based approaches. Finally, we summarize the major challenges that remain and outline promising avenues for future research.
keywords:
Magnetic Resonance Imaging , Image Harmonization , Deep Learning
[…]
booktabs
1 Introduction
Figure 1: Overview of harmonization strategies spanning the entire medical imaging pipeline, including image acquisition, reconstruction, post-processing, and feature-level analysis. This figure shows representative examples of image contrasts and image-based measurements, such as DWI: diffusion weighted image; T1WI: T1 weighted image; T2WI: T2 weighted image; RBV: Regional brain volume; CT: Cortical thickness; FA: Fractional Anisotropy.
Magnetic resonance imaging (MRI) has had a profound impact on the field of medicine, with widespread applications in medical and neuroscience research, computer-aided diagnosis, longitudinal monitoring, and image-guided interventions. To advance scientific discovery and bridge the gap between research and clinical practice, collecting and sharing large-scale imaging datasets across sites has become increasingly essential [123, 120, 81, 111, 107, 17]. Multi-center studies that aggregate large and diverse samples not only enhance statistical power, particularly important for investigating rare or low-prevalence diseases, but also provide broader coverage of key biological variables such as age, sex, race, geographic location, socioeconomic status, and disease subtypes. The increased sample size and heterogeneity also improve the ability of studies to detect subtle yet meaningful effects in high-dimensional spaces of variables and confounders [82, 13, 13].
In the era deep learning, the proper collection and analysis of large-scale medical imaging data has become even more critical [136, 75]. Although machine or deep learning models have shown great potential in addressing scientific challenges in medicine, their applicability in clinical practice remains limited. Many of these models are built on the assumption that the training and testing datasets come from the same distribution, and their performance can degrade substantially when this assumption does not hold. In other words, when applied to images acquired using protocols that are different from those used during training, these models often exhibit poor reproducibility and limited generalizability [46]. As a result, their effectiveness may decline significantly when used across different hospitals, scanners, or patient populations. This decline in performance is largely caused by non-biological technical variability, often referred to as scanner effects, device effects, or batch effects, which can stem from differences in scanner/device hardware and software by different manufacturers, sequences, acquisition protocols, image reconstruction pipelines and techniques, and other sources [79, 22, 59, 48].
A common conventional approach to avoid the challenges of directly using and comparing heterogeneous imaging data is meta-analysis, where each site performs its own analysis independently, and the results are later combined [3, 68, 30]. However, meta-analyses typically include only group-level statistical and clinical information, making it difficult to perform detailed modeling or adjustments at the individual level. Moreover, when participant distributions are imbalanced, site-specific statistics may introduce systematic biases. In studies with small imaging sample sizes, fluctuations in parameter estimation during z-score conversion may further compromise the stability of statistical inference. In contrast, mega-analysis involves the joint analysis of all raw imaging data on a unified platform [82, 78, 23, 57, 53]. However, this strategy places more stringent demands on data harmonization, as combining datasets from different centers may introduce additional non-biological variability, particularly due to differences in imaging protocols. Therefore, effective harmonization is a critical prerequisite for the success of mega-analyses.
Comprehensive MRI harmonization involves three key components: harmonized acquisition, image-level processing, and feature-level analysis. Harmonized acquisition refers to the prospective standardization of the use of scanner hardware, pulse sequences, and protocol parameters during data collection, aiming to reduce variability or heterogeneity of the acquired data [71]. Image-level harmonization involves retrospective adjustments to the acquired images, such as intensity normalization, statistical correction, or deep learning methods to standardize image contrast and signal distribution [57, 1, 31]. Feature-level harmonization focuses on quantitative metrics, texture features, and anatomical representations extracted from images, ensuring their comparability across different sites to support reliable data integration and cross-site analysis [39, 135, 38]. The primary goal of harmonization is not to recover some absolute “ground truth,” but rather to enhance the reliability of comparisons at both individual and group levels. In other words, the essence of harmonization is to eliminate non-biological technical variation while preserving meaningful biological differences. Rather than removing systematic bias, harmonization seeks to make such biases consistent across all datasets, thereby minimizing their impact on downstream analyses.
This review focuses on multi-site harmonization methods for MRI data, although the underlying principles are broadly applicable to other medical imaging modalities. Several review papers have been published on this topic, covering modalities such as positron emission tomography, computed tomography, and microscopic pathology [57, 1, 91, 126, 9, 87]. However, the primary emphasis remains on MRI, owing to its inherent characteristics, such as multiple field strengths, diverse imaging modalities, and a wide range of quantitative parameters, which result in pronounced inter-site heterogeneity. Moreover, compared to other modalities, MRI is uniquely complex due to the breadth of imaging capabilities it offers, presenting a wider range of harmonization challenges and thus requiring more comprehensive solutions. Nevertheless, existing MRI harmonization reviews have largely focused on retrospective approaches applied to structural and diffusion MRI [57, 1, 91], while prospective strategies at the acquisition stage have received limited attention. In this review, we place particular emphasis on recent advances in vendor-agnostic pulse sequences and harmonized image reconstruction techniques. For retrospective harmonization, we highlight emerging deep learning-based methods that have gained traction in recent years. This is the first comprehensive review to cover the full spectrum of harmonization efforts across the MRI pipeline - from prospective acquisition harmonization to retrospective image- and feature-level methods Figure 1.
2 Background
Table 1: Confounding factors in medical imaging, in specific in MRI, data
Measurement Bias (non-biological) Sampling Bias (biological)
Scanner Properties (hardware): - Scanner manufacturer - Field strength - Gradient Systems - Radiofrequency Coils - Hardware Imperfections - Field inhomogeneity ($\Delta B_{0}$, $B_{1}^{+}$, $B_{1}^{-}$) Biographic Data: - Age - Age by disease appears - Gender - Background population (e.g., ethnicity, genetics)
Sequence (software): - Sensitivity to tissue differences - Sensitivity to motion and flow - Vendor-specific underlying implementations Severity of a Condition of Interest (e.g., psychiatric disease)
Acquisition Parameters (software): - Spatial resolution: field of view (FOV), matrix size, slice thickness, - Signal-to-noise ratio (SNR): flip angle, receiver bandwidth, number of excitations - Contrasts: echo time (TE), repetition time (TR), inversion time (TI), magnetization preparation/suppression, b-value, number of gradient directions Study Protocol: - Healthy control inclusion criteria - Case inclusion criteria - Diagnostic criteria and instrument
Image Postprocessing (software): - Image normalization and filtering - Image reconstruction and multi-coil combination - Distortion, eddy current and other imperfections correction Neurobiological and Medical Comorbidities: - Past medical history - Current medical conditions - Past and current medication
Site differences are even more heterogeneous in diffusion and functional MRI due to echo planar imaging (EPI) -induced effects such as phase differences, image distortions and signal dropouts or detailed experimental setup such as variable acquisition geometry, brain coverage, difference in task implementation and temporal duration.
2.1 Confounding factors in MRI data
The collection and analysis of medical imaging data involve a complex and variable set of procedures, including subject recruitment/selection, imaging hardware and protocol design, and the choice of data analysis models. Variations at any of these stages (across imaging sites, subjects, or even repeated scans) can introduce differences into the final results [33, 7, 95]. Therefore, to achieve effective data harmonization, i.e., the removal of non-biological variability while preserving biologically meaningful differences across participants and sites, it is essential to understand the sources of such variability. The major contributing factors can be categorized into measurement (non-biological) bias and sampling (biological) bias, as summarized in Table 1 for magnetic resonance imaging.
2.2 Prospective and retrospective approaches
Prospective harmonization refers to strategies that are deliberately planned prior to data acquisition with the goal of minimizing anticipated sources of variability at the source [23, 57, 71]. A common approach involves standardizing scanner models and acquisition protocols in advance to reduce differences introduced during image acquisition. Building on this foundation, this review highlights recent advances such as vendor-agnostic pulse sequences and harmonized image reconstruction methods, which aim to overcome traditional barriers to acquisition consistency through open-source and easily implementable solutions. These innovations further enhance the effectiveness of prospective harmonization. In addition, the use of traveling subjects represents another important prospective strategy, whereby the same individuals are scanned across multiple sites to obtain matched datasets. This design provides a valuable reference for quantifying and correcting systematic inter-site differences during analysis, thereby supporting the effective removal of non-biological variability.
Retrospective harmonization refers to the process of applying harmonization techniques after data acquisition to existing, heterogeneous multi-site datasets, with the aim of eliminating or significantly reducing variability caused by non-biological factors [57, 91, 84, 65]. These methods currently dominate the field, largely due to the accessibility of large-scale public multi-site datasets and their advantages in terms of cost-effectiveness and flexibility compared to prospective approaches. Current retrospective strategies encompass conventional image processing, statistical modeling, and deep learning-based methods. Their effectiveness largely depends on the ability to accurately identify, model, and account for the sources, structure, and manifestations of batch/site effects in the data, as well as the precision with which the algorithm can distinguish true biological signals from technical noise.
2.3 Image-level and feature-level approaches
Image-level harmonization aims to directly modify the voxel intensities of MRI scans and is typically framed as an image-to-image translation task. The goal is to make images acquired from different sources be similar in terms of characteristics such as contrast, sharpness, and signal-to-noise ratio (SNR), as if they were acquired under comparable settings. In principle, harmonized images can then be used for a variety of downstream tasks such as segmentation and feature extraction, and may also be visually inspected by radiologists. However, image-level harmonization carries the risk of altering underlying anatomical structures or introducing artifacts, especially when using complex generative models based on deep learning. Therefore, it is crucial to properly validate techniques by validating the biological fidelity of the harmonized images [57, 31, 126, 87].
Feature-level harmonization methods operate on quantitative features or image-derived measurements extracted from raw MRI data, such as regional brain volumes, cortical thickness, diffusion tensor imaging (DTI) metrics, functional connectivity matrices, or radiomic features. This makes them well-suited for incorporating statistical models to remove site effects. One key advantage of feature-level approaches is their ability to directly include biological covariates in the modeling process. When the feature extraction is robust, these methods typically pose a lower risk of altering image appearance or anatomical structures, and they are generally more computationally efficient than image-level deep learning methods. However, because they rely on image-derived features, the quality and relevance of the selected features have a direct impact on the effectiveness of harmonization. Moreover, harmonization is restricted to the specific features extracted and may not generalize to other types of analysis performed on the same images [57, 39, 5, 58].
3 Harmonized Data Acquisition
A medical imaging device consists of two main components: hardware and software (Figure 1). For an optimally harmonized acquisition across multiple sites, it is best if all the components of the hardware and software could be exactly matched. In MRI, for example, this would mean scanners of the same type with the same (main and gradient) field strengths, and the same receive and transmit coils are used. However, this may pose significant restrictions on large-scale studies especially in global health studies where access to the best devices that are often needed for cutting-edge research (such as connectome strength magnets), may be limited. While matching device hardware may not always be feasible, often there is more flexibility around the software components that are used at the acquisition level, which is discussed in the following subsections. Software could also be used in post-acquisition processing to compensate for the device hardware differences, that are discussed in Section 4 and Section 5.
3.1 Vendor-agnostic pulse sequence
Pulse sequences serve as the core of MR image formation; thus, their harmonization offers a principled approach to addressing site effects at the source. However, due to differences in the underlying implementation of pulse sequences from different vendors, signal discrepancies may still arise even when identical acquisition parameters listed in Table 1 (e.g., TE, TR, FOV, matrix size) are used [64]. Such variations originate from factors including, but not limited to, differences in fat suppression modules, dephasing strategies (e.g., spoiler and crusher gradients), RF pulse shapes and profiles, most of which are not accessible or adjustable through the user interface [71, 41]. To address this challenge and enhance consistency at sequence level, several vendor-agnostic or open-source pulse sequence platforms have been developed over the past decade, including Pulseq [71], gammaSTAR [70] and RTHawk[99]. For example, Pulseq enables modular pulse sequence programming in MATLAB and Python, allowing extensive and detailed control over RF pulses, gradient waveforms, and inter-module interval. The resulting sequence is compiled into a standardized .seq file, which can then be interpreted and executed by vendor-specific backends for MRI scanning (Figure 2). Additionally, Pulseq can be integrated with various MRI simulation and graphical sequence design tools, further alleviating the steep learning curve of pulse sequence development.
Figure 2: Harmonized acquisition and reconstruction workflow proposed in [41]. The pulse sequence was implemented using Pulseq to ensure identical configurations across scanners and vendors. All post-processing steps including image reconstruction and quantitative parameter fitting were performed offline using a consistent pipeline.
Liu et al. [76] systematically evaluated a single-shell diffusion MRI sequence implemented using Pulseq across two scanners from different vendors, using standard error as a metric of repeatability. For mean diffusivity in phantoms, the Pulseq sequence demonstrated 2.5-fold superior inter-scanner reproducibility compared to vendor-provided sequences. In human brain imaging, a Pulseq sequence reduced inter-scanner standard error in fractional anisotropy by 35–50% across various brain regions. In addition to diffusion MRI, vendor-agnostic pulse sequence tools have also been validated in quantitative MRI applications, including chemical exchange saturation transfer [54], brain T1 and T2 mapping [41, 67], and myocardial T1 mapping [44, 43]. In the work by Karakuzu et al. [64], combining RTHawk-based acquisition harmonization with a unified parameter quantification workflow led to statistically significant reductions in inter-vendor variability for T1, magnetization transfer ratio, and magnetization transfer saturation index measurements.
Although current preliminary results are encouraging, it remains unclear to what extent vendor-agnostic or open-source pulse sequence tools can mitigate site effects, as they have yet to be widely implemented or validated at scale. On one hand, while current frameworks allow for setting basic hardware constraints or upper limits, they may not sufficiently account for hardware-specific differences, which can continue to contribute to inter-site variability. On the other hand, in the absence of direct vendor support and integration, the implementation and validation of new sequences, as well as obtaining regulatory clearance for research or clinical use, can be time-consuming and resource-intensive. For example, due to the need to bypass vendor-specific post-processing systems, designing raw data processing pipelines often demands substantial technical expertise and effort, and may still rely on customized reconstruction workflows [114].
3.2 Harmonized data reconstruction
The raw MRI signal acquired from the scanner is one-dimensional complex-valued data. To generate the final image, this signal should be filled into a predefined k-space trajectory and then transformed via Fourier transformation, which is known as image reconstruction. Differences in reconstruction pipelines can introduce non-negligible variability, contributing to a lack of harmonization across sites or vendors. These differences may arise from multiple factors, including pre- and post-reconstruction distortion and phase correction, k-space gridding, partial Fourier reconstruction, multi-coil parallel reconstruction, and coil combination strategies [49]. While vendors provide access and control over some of the options, filters, and parameters through user interfaces during acquisitions, a full control over the entire raw data processing pipeline often requires additional programming within the vendor software environment, which may not always be available or if available, may not be straightforward.
Offline open-source reconstruction toolboxes offer alternative, promising opportunities for standardizing the image reconstruction process. Representative examples include the Berkeley Advanced Reconstruction Toolbox (BART) [15, 83] and the Michigan Image Reconstruction Toolbox (MIRT) [37]. For instance, BART not only implements conventional parallel imaging algorithms but also provides general-purpose solutions for non-Cartesian, model-based, and deep learning-based reconstruction. Its cross-platform, open-source, and multi-language support (Linux terminal, MATLAB, and Python) make it accessible and easy to integrate into diverse research workflows. Prior to reconstruction, inconsistencies in raw data formats across vendors pose a significant challenge. The ISMRMRD (International Society for Magnetic Resonance in Medicine Raw Data) [61] framework addresses this issue by providing a standardized format that harmonizes vendor-specific raw data and headers, thereby facilitating consistent and reproducible reconstruction pipelines.
Despite their ease of use, offline open-source reconstruction toolboxes have inherent limitations that restrict their clinical scalability. These include the lack of real-time quality control and the requirement to store large raw datasets. To address these challenges, online reconstruction frameworks such as Gadgetron [50] have been proposed. Gadgetron adopts a modular, streaming-based architecture and incorporates a wide range of extensible toolboxes, enabling real-time reconstruction through GPU or multithreaded CPU acceleration. It also supports advanced features such as automated motion tracking and scan planning. A notable example is the HERON framework developed by Jordina et al. [121], which applies online reconstruction to motion-sensitive fetal diffusion MRI (dMRI). This approach leverages image-based real-time motion estimation to dynamically adjust fetal dMRI acquisition, thereby mitigating the impact of unpredictable fetal motion. In this context, Gadgetron is used to enable real-time landmark detection through deep learning and motion analysis, providing dynamic feedback for prospective adjustment of the acquisition. While fetal MRI remains one of the challenging applications of MRI due to the non-periodic, non-rigid, and complex fetal and maternal body movements, similar applications have been developed for fetal functional MRI [104], fetal cardiovascular MRI [105], and quantitative MRI of the fetal brain and placenta [122]. Online reconstruction in these challenging applications enables motion-informed quantification and adaptive data re-acquisition, further promoting data consistency across scans and sites.
4 Image-level Retrospective Harmonization
4.1 Traditional image-level approaches
Traditional image-level harmonization methods primarily rely on various intensity normalization techniques. Although these methods are not explicitly designed to remove site effects, they are commonly used as preprocessing steps or baseline approaches due to their simplicity and low computational cost. Such methods typically apply global transformations to the entire image (e.g., z-score normalization), adjust image intensity statistics (e.g., histogram matching), or utilize reference intensities from specific tissue types to align global or local intensity distributions, thereby improving comparability across scans acquired from different sites.
A widely used class of methods is based on histogram matching [89, 100], which aims to align the intensity histogram or cumulative distribution function (CDF) of a source image with that of a target image or a predefined reference distribution. While these methods are conceptually simple and computationally efficient, they are often sensitive to outliers (e.g., hyperintense lesions) and may fail to preserve biologically meaningful variations at the individual level. Another class of methods relies on reference-based normalization, such as White Stripe proposed by Shinohara et al. [102], which rescales image intensities using a reference region composed of normal-appearing white matter. Building upon this approach, RAVEL (Removal of Artificial Voxel Effect by Linear regression) [40] further addresses residual non-biological variability that may persist after White Stripe normalization, which will be introduced in Section 5.1.
For diffusion MRI data, one approach that performs harmonization on the raw acquisition signal prior to feature extraction is the RISH (Rotationally Invariant Spherical Harmonics) [84, 65] method, which can serve as a preprocessing step compatible with subsequent analysis pipelines. RISH represents voxel-wise diffusion signals as projections onto a set of orthogonal basis functions defined on the unit sphere. Due to its rotational invariance, RISH features are robust to variations in gradient directions across scans. Although RISH has been shown to effectively preserve biological effects, it relies on feature-matched healthy controls or traveling subjects across sites for calibration. To address this limitation, De Luca et al. [29] proposed modeling RISH features using a covariate-driven general linear model (RISH-GLM), which enables multivariate modeling and cross-site harmonization without requiring matched training data.
4.2 Learning-based image-level approaches
Learning-based image-level harmonization methods are predominantly driven by deep convolutional neural networks and can be broadly categorized into four groups: adversarial learning and style transfer, anatomy-contrast disentanglement, multi-contrast prior learning, and source-free distribution modeling. A general trend across these approaches is the shift from fixed, site-specific harmonization toward adaptive multi-site solutions, and from methods requiring simultaneous access to data from multiple sites to those leveraging only single-site data.
4.2.1 Adversarial learning and style transfer
For learning-based approaches, Image-level harmonization aligns with the fundamental characteristics of conventional image style transfer. Consequently, generative adversarial networks (GANs) and their variants, as leading techniques in style transfer tasks, were among the first methods employed to address the challenge of unpaired image harmonization [137] (Figure 3a), with Cycle-consistent GAN (CycleGAN) being the most widely used approach [85, 113, 20, 93]. The basic strategy involves treating unpaired data from two different sites as the source and target domains, respectively, and applying cycle-consistency loss to transfer style information while preserving biological structures. Building on this foundation, improvements to the generator/discriminator architecture, modifications to loss functions, or the incorporation of attention mechanisms have been explored to enhance training stability and harmonization performance.
Figure 3: Four representative categories of image-level deep learning-based harmonization methods: (a) adversarial learning and style transfer, (b) anatomy-contrast disentanglement, (c) multi-contrast prior learning and (d) source-free distribution modeling. A: anatomy; S: structure; C: content.
A major limitation of CycleGAN is that it requires separate training for each pair of sites and fails to leverage complementary information across multiple sites. To address this, several many-to-one GAN-based harmonization strategies have been proposed [97]. For example, the IGUANe framework proposed by Roca et al. [97] employs a single universal generator capable of translating images from multiple source sites into the style of a reference site. In practice, this universal generator is adversarially trained against multiple site-specific discriminators, while multiple backward generators are trained similarly to enforce cycle consistency. In a harmonization task involving 11 sites/scanners, the IGUANe architecture required training 11 generators and 20 discriminators, resulting in substantial computational overhead. StyleGAN [66] and StarGAN [24, 25] further extend CycleGAN by introducing explicit site or style encodings, which significantly reduce the number of required encoders and discriminators. For instance, Bashyam et al. [8] achieved harmonization across six sites using only four models based on the StarGAN framework, and demonstrated improved accuracy in brain age prediction in downstream tasks.
4.2.2 Anatomy-contrast disentanglement
GAN-based image style transfer offers a straightforward solution for harmonization, but it lacks explicit separation between anatomical structures and contrast-related features. In contrast, variational autoencoder (VAE), which encode images into low-dimensional latent representations and decode them back to images, can be used to explicitly disentangle anatomy and contrast (Figure 3b). In the MURD method proposed by Liu et al. [77], independent style encoders, content encoders, and image generators are introduced. However, training such a framework with unpaired multi-site data presents significant challenges. To address this, MURD additionally incorporates a style generator and a discriminator, and constrains the learning process using as many as seven loss components, including adversarial loss, cycle-consistency loss, and content/style consistency losses. Once the latent features representing anatomy and style are explicitly disentangled, image harmonization can be achieved by recombining the anatomical representation of the source domain with the style of the target domain. This strategy not only accommodates resolution differences across sites, but also enables continuous harmonization of imaging styles between sites.
Although the use of multiple loss functions and adversarial training has proven effective, such training strategies are often unstable and require careful tuning of the weighting coefficients for each loss term. In contrast, ImUnity, proposed by Cackowski et al. [18], adopts a contrastive learning strategy to simplify the training process. Specifically, a VAE is first trained in a self-supervised manner to generate realistic images. For each subject, two 2D slices from different anatomical locations are randomly selected: one slice remains unaltered to extract anatomical information, while the other undergoes a random gamma transformation. The VAE is then trained to generate a stylized image that retains the anatomy of the first slice but adopts the style of the second. Images from multiple sites are used to train this model, with the objective of removing site-specific information while preserving biological content. At inference time, harmonization is achieved simply by feeding the source image and a reference image into the model, without requiring fine-tuning on unseen sites. Unlike ImUnity, DLEST, proposed by Wu et al. [129], performs style transfer directly in the latent space. First, site-invariant image generation is trained using single-site data to obtain an encoder for extracting latent representations and a decoder for reconstructing images. Then, latent representations from different sites are introduced into the generator, and an energy-based model is trained to transform source-site latent codes into those of the target site. The harmonized image is then generated by decoding the transformed latent representation using the trained decoder.
4.2.3 Multi-contrast prior learning
Clinical acquisitions and large-scale public MRI datasets often include multi-contrast images, such as T1, T2, PD (proton density) and FLAIR (fluid attenuated inversion recovery), to capture complementary tissue properties. It is commonly assumed that these contrasts, when acquired from the same subject, share a consistent anatomical structure, thereby providing naturally paired anatomy-contrast information for disentanglement (Figure 3c). This mitigates concerns regarding pathology or demographic differences between the groups and reduces, or even eliminates, the need for simultaneous access to both source and target datasets.
Dewey et al.[32] used co-registered T1- and T2-weighted images from the same subject as input to the encoder to disentangle the latent space and obtain anatomical one-hot encoding maps ($\beta$) and contrast encoding vectors ($\theta$). During training, components were randomly sampled from $\beta$ and $\theta$, recombined, and passed to the decoder, with self-supervised losses enforcing intra-subject consistency in $\beta$ and protocol-dependent variation in $\theta$ to ensure anatomy-contrast disentanglement. During inference, images of arbitrary contrast from a new site can be encoded to obtain $\beta$, which is then combined with a fixed $\theta$ from a reference site and decoded to produce harmonized images. The CALAMITI method, proposed by the same group, builds upon Zuo et al. [139] by introducing cross-contrast synthesis and adversarial learning across different sites. These enhancements enable CALAMITI to achieve more robust disentanglement and to learn a globally consistent anatomical representation across sites. In addition, CALAMITI incorporates a 3D fusion network, making it applicable to volumetric data and demonstrating improved performance in downstream segmentation tasks. Furthermore, the HACA3 method proposed by Zuo et al. [140] incorporates more than two MRI contrasts (T1, T2, PD and FLAIR) and supports flexible combinations of available modalities, making it well-suited for clinical scenarios where multi-contrast data may be incomplete or heterogeneous. Compared to CALAMITI, which assumes full structural consistency across contrasts, HACA3 introduces contrast- and artifact-aware attention to address subtle anatomical differences and artifact contamination in multi-contrast MRI, enabling more robust and flexible harmonization.
In addition to the structural similarity, multi-contrast MR images are linked through inherent physical properties. Specifically, different contrast-weighted images acquired via varying pulse sequences and acquisition parameters ultimately reflect the same underlying tissue properties, i.e. the “quantitative” relaxation times T1 and T2, and PD. Based on this, Qiu et al. [94] and Borges et al. [16] proposed physics-driven harmonization frameworks that translate multi-contrast MRI into quantitative maps that are invariant to modality and acquisition protocol. In these approaches, contrast-specific forward physical models are embedded into the loss function to enable self-supervised learning. Additionally, PhyCHarm, proposed by Lee et al. [72], further leverages scanner-specific acquisition parameters to synthesize native MR images from the quantitative maps for harmonization. However, the final step still requires supervision from paired training data.
4.2.4 Source-free distribution modeling
To eliminate the dependency on multi-contrast and multi-site data while ensuring generalizability to unseen domains, a new generative model, normalizing flow, was introduced to directly model the source distribution. This enables source-free harmonization without the need for traveling subjects, multi-contrast data, or task-specific supervision, as the model learns to map a complex probability distribution to a simple one through a series of invertible and differentiable transformations.
Jeong et al. and Beizaee et al. independently and almost simultaneously introduced normalizing flows into the field of image harmonization, proposing the BlindHarmony [62] and Harmonizing flows [11] frameworks, respectively. Taking Harmonizing flows as a representative example, the method adopts a three-step harmonization strategy consisting of source domain modeling, harmonizer pre-training, and test-time domain adaptation. Specifically, a normalizing flow network, composed of stacked affine coupling layers, was first trained to capture the distribution of source domain images. This model provides an invertible and differentiable transformation that maps the complex source distribution to a standard Gaussian. Subsequently, a lightweight UNet was employed as the harmonizer and pre-trained to reconstruct original source images from their randomly augmented counterparts, thereby learning to compensate for appearance variations introduced by contrast and intensity shifts. Finally, during inference, the harmonizer was fine-tuned using target domain images under the supervision of the frozen flow model, ensuring that the harmonized outputs align with the source domain distribution (Figure 3d). This framework achieves unsupervised, source-free and task-agnostic harmonization, and demonstrates generalizability to unseen domains, with robust performance across multiple tasks, including brain MRI segmentation and neonatal age estimation.
5 Feature-level Retrospective Harmonization
5.1 Statistical approaches
Feature-level methods based on statistical modeling typically assume that extracted features (e.g., brain volumes, cortical thickness, DTI metrics) can be decomposed into biological effects, site or batch effects, and random noise. By fitting a statistical model (most commonly a linear model), the site effect can be estimated and subsequently removed or adjusted, yielding harmonized data. A comprehensive review of such methods is available in Hu et al. [57] Here, we briefly introduce several representative approaches, with a particular emphasis on ComBat [39, 63], which serves as a foundation for subsequent learning-based methods.
The ComBat model was originally developed for batch effect correction in gene expression data [63], and was first introduced to DTI data by Fortin et al. [39] In this model, the observed feature $y_{ijv}$ at voxel $v$ for subject $j$ from site $i$ is modeled as a linear combination of multiple factors. These factors include the global mean $\alpha_{v}$, biological covariates (e.g., age and sex), and site effects (additive and multiplicative effects, $\gamma_{iv}$ and $\delta_{iv}$). The full model can therefore be expressed as:
$y_{ijv}=a_{v}+\mathbf{X}_{ij}\beta_{v}+\gamma_{iv}+\delta_{iv}\varepsilon_{ijv}$ (1)
where $\mathbf{X}$ is the design matrix for the covariates, $\beta$ represents the corresponding regression coefficients, and $\varepsilon$ is the error term, assumed to have zero mean and variance $\sigma^{2}$. After estimating the coefficients using the empirical Bayes method, the ComBat-harmonized value can be expressed as:
$y_{ijv}^{\mathrm{ComBat}}=\frac{y_{ijv}-\hat{a}_{v}-\mathbf{X}_{ij}\hat{\beta}_{v}-\hat{\gamma}_{iv}}{\hat{\delta}_{iv}}+\hat{a}_{v}+\mathbf{X}_{ij}\hat{\beta}_{v}$ (2)
Based on this, ComBat has been shown to be effective across a variety of imaging features, including not only DTI-derived metrics but also cortical thickness, functional connectivity spectroscopy and radiomics [135, 38, 95, 127, 2, 12, 69]. In addition, numerous extensions of the standard ComBat model have been proposed to relax the assumptions of the original version, including the use of alternative parameter estimation strategies and applications to more complex study designs [92, 56, 96, 117, 19, 27, 138, 131]. However, one major limitation of this method is that it requires each scanner to have a statistically representative sample. This constraint reduces its generalizability to unseen data.
Unlike ComBat, which explicitly models additive and multiplicative effects, some other strategies model biological factors or site effects using basis representation or latent factors [40, 36, 21, 73, 98]. Taking RAVEL [40] as an example, it firstly selects a control voxel that is highly sensitive to variations in reconstruction algorithms, acquisition protocols, and scanner configurations, typically from the cerebrospinal fluid (CSF), to serve as a proxy for non-biological effects. Then, RAVEL performs singular value decomposition on the control voxels to extract latent factors representing technical variation, and then applies linear regression across all voxels to estimate and remove these effects. Finally, the resulting residuals are treated as the RAVEL-corrected intensities.
5.2 Learning-based feature-level approaches
Learning-based feature-level approaches are typically extensions of statistical harmonization strategies, particularly those derived from ComBat. A representative example is Neuroharmony [42, 6], which is based on the assumption that the intrinsic image characteristics of a single image can aid in data harmonization. This approach addresses the limitation of traditional ComBat, which cannot generalize to unseen sites. Specifically, Neuroharmony first applies ComBat to existing multi-site data to obtain corrected features for each image. Then, using the MRIQC tool, a range of image quality metrics (IQMs) are extracted from each image, including SNR, contrast, blurriness, motion artifacts, and background uniformity. Based on these metrics, Neuroharmony employs a random forest model to learn the mapping between the 64 IQMs and the ComBat-derived corrected features. Once trained, the model no longer relies on population-level statistical features but instead performs harmonization using only the image’s IQMs and biological covariates.
Another line of learning-based feature-level approaches leverages conditional variational autoencoder (cVAE) [86] to address nonlinear site-related variations and support multivariate modeling. In the cVAE framework, an encoder first processes feature vectors (e.g., ROI-based cortical thickness or SH representations) to generate latent representations. These representations are then concatenated with site information (i.e., a one-hot vector) or biological covariates and fed into a decoder to reconstruct the original feature vectors. To accommodate 1D input, both the encoder and decoder are typically implemented as fully-connected neural networks. To encourage site-invariant latent representations, mutual information between the latent features and the site encodings is minimized during training. Then, in the harmonization phase, modifying the input site encoding allows for flexible translation of input features to any target site.
Several extensions of the cVAE framework have been developed to enhance harmonization performance. For instance, the goal-specific cVAE (gcVAE) [4] proposed by An et al. incorporates a pretrained classifier into the standard cVAE architecture. This allows the original cVAE to implicitly preserve biologically meaningful representations by leveraging supervision from downstream classification tasks during training. Another variant, DeepComBat by Hu et al. [58], integrates cVAE with the classical ComBat method. It first applies ComBat to the latent mean vectors produced by the cVAE encoder, followed by decoding the harmonized latent representations to reconstruct the original features. A second ComBat step is then applied to the residuals (i.e., the difference between the reconstructed and original features) to remove residual site effects, which are subsequently added back to produce the final harmonized output.
Distribution differences in covariates (e.g., age and sex) are common and often unavoidable in multi-site datasets. As theoretically demonstrated by Tachet et al. [26], directly applying cVAE under such conditions may lead to covariate-driven variations being incorrectly attributed to site effects. To address this issue, DeepResBat, proposed by An et al. [5], introduces a two-stage strategy that performs feature harmonization after explicitly accounting for covariate effects. Specifically, it first estimates covariate influences using nonlinear regression models. The covariate residuals, obtained by subtracting the estimated covariate contributions from the original features, are then used as input to a cVAE model to isolate and remove site-specific effects, yielding harmonized residuals. The final harmonized features can be reconstructed by reintroducing the covariate effects into the harmonized residuals (Figure 4). By targeting residuals rather than raw features for deep learning harmonization, DeepResBat explicitly preserves biological variability while effectively reduces the risk of spurious associations.
Figure 4: Model structure for the feature-level deep learning harmonization method DeepResBat, proposed by [5]. The covariates effect of the original features were first removed by subtraction and used as the input to a VAE. Then, the VAE output was added back to the removed covariate components to obtain harmonized features. ROI: region of interest.
6 Traveling Subject
Harmonization models based on non-traveling subject datasets, whether traditional statistical approaches or learning-based methods, can effectively eliminate site effects. However, it remains unclear whether such models may also overcorrect biological variability across sites. In this context, traveling subject-based approaches offer a baseline for rigorously disentangling biological and non-biological sources of variability. Moreover, publicly available traveling subject datasets not only enable investigation into how site effects influence multi-site statistical analyses, but also serve as valuable benchmarks for validating newly proposed harmonization methods.
6.1 Statistical approaches
Building on image-based histogram matching methods, Wrobel et al. proposed Multisite Image Harmonization by Cumulative Distribution Function Alignment (MICA) [128], which performs image harmonization based on the alignment of CDFs. The method first applies preprocessing steps such as N4 bias field correction and skull stripping to the images, and then computes their CDFs. For each traveling subject, MICA selects one image as the template and uses its CDF as the alignment target. It then estimates a nonlinear, monotonically increasing warping function by densely sampling paired points between the source and template images and applying linear interpolation. This warping function is used to align the CDF of the source image to that of the template.
Based on traveling subject (TS) data, ComBat can also be extended to the TS-ComBat method [80]. In this approach, the covariate term in the standard ComBat model is replaced with individual effects estimated from traveling subjects, while the site effects are still estimated and removed using an empirical Bayes method. To further account for repeated measurements across time points, Beer et al. [10] applied Longitudinal ComBat in the context of traveling subject studies. The model is expressed as follows:
$y_{ijv}(t)=a_{v}+\mathbf{X}_{ij}(t)\beta_{v}+\eta_{jv}+\gamma_{iv}+\delta_{iv}\varepsilon_{ijv}(t)$ (3)
where both $y_{ijv}(t)$ and $\varepsilon_{ijv}(t)$ introduced time-varying dependent variables, $\eta_{jv}$ represents subject-specific random intercept. In addition, several of the previously discussed image-level and feature-level harmonization methods have been further validated and extended on traveling subject datasets, demonstrating their adaptability across sites [80, 28, 132].
6.2 Learning-based approaches
Traveling-subject data naturally provide paired training samples for learning-based methods, offering a more direct solution compared to unpaired approaches. Methodologically, these techniques can be categorized into two main types: end-to-end mapping and anatomy-contrast disentanglement strategies similar to those used in unpaired settings (Figure 5).
Figure 5: Two representative deep learning-based harmonization strategies using traveling subject data: (a) end-to-end mapping and (b) anatomy-contrast disentanglement methods. The availability of paired training data provides additional supervision related to site or subject identity, which enhances the learning of site-invariant representations. A: anatomy; S: structure.
6.2.1 End-to-end mapping
Based on paired training data, Dewey et al. proposed DeepHarmony [31], a U-Net-based harmonization framework. In their study, 12 subjects were scanned on two different scanners using protocols (e.g., T1, T2, PD, and FLAIR) with varying parameters. Prior to end-to-end learning, extensive preprocessing steps were applied to the multi-scanner data, including bias field correction, resolution enhancement, image registration, and gain correction. Three separate U-Nets were trained on axial, sagittal, and coronal slices, respectively, and final 3D volume reconstruction was achieved by voxel-wise median fusion across the three orientations. Finally, DeepHarmony was shown to substantially reduce inter-protocol volumetric discrepancies in longitudinal MRI datasets of patients with multiple sclerosis.
Unlike methods that perform direct mapping in the image domain, Tong et al. [115] proposed a harmonization approach that maps source diffusion-weighted images to target diffusion kurtosis imaging (DKI) parameters. A 3D hierarchical convolutional neural network was trained using co-registered labels estimated through an iteratively reweighted linear least squares method. This approach resulted in a 50-60% reduction in inter-scanner variation of DKI parameters within white matter. Similarly, Tax et al. [110] and Ning et al. [88] summarized several learning-based harmonization methods from the Multi-Shell Diffusion MRI Harmonization Challenge (MUSHAC). These methods harmonize DWI data in the spherical harmonics domain and include: a basic fully connected network, approaches incorporating residual learning and spherical convolutions to improve training efficiency and reconstruction accuracy, a fully convolutional shuffle network and a sparse dictionary learning-based method. All included methods significantly reduced variability across multi-scanner DWI acquisitions, although challenges remain in accurately capturing localized features.
To enhance the generalizability of harmonization frameworks to data from unseen sites, Xu et al. proposed Site Mix (SiMix) [130], a method based on mixed training data and test-time perturbation. Rather than selecting one existing site as the harmonization target, SiMix generates a virtual target site for model training by computing a weighted combination of images from multiple known sites. Once the model is trained, the initially harmonized virtual-site image is then linearly mixed with the original test image to produce multiple mixed test images, which are then passed through the trained model again. Following the ensemble learning strategy, the final harmonized result is obtained by averaging the outputs across all mixed inputs, thereby improving adaptability to arbitrary test domains.
6.2.2 Anatomy-contrast disentanglement
Compared to non-traveling-subject methods, traveling-subject-based anatomy-contrast disentanglement offers the distinct advantage of efficiently utilizing shared anatomical structures across different sites for strong supervision, thereby enabling more accurate interpretation and quantification of site and scanner effects (Figure 5b).
A representative method is Multi-scanner Image harmonization via Structure Preserving Embedding Learning (MISPEL), proposed by Torbati et al. [118] The framework consists of scanner-specific encoders and decoders and follows a two-stage training strategy. In the embedding learning stage, a 2D U-Net serves as the encoder to extract latent embeddings (multiple 2D feature maps) from the source images, while the decoder reconstructs the target image through a linear combination of these embeddings. Slices with matched anatomical content from different scanners are used to jointly train the encoder-decoder, with the loss computed to enforce consistency of the latent embeddings across scanners by minimizing their pixel-wise variance. In the harmonization learning stage, the encoder is frozen and only the decoder is trained, aiming to minimize the difference between reconstructed images of the same slice from different scanners while preserving fidelity to the original image. This approach demonstrates that paired multi-site data can provide strong supervision, enabling the model to maintain high anatomical fidelity during harmonization. Notably, ESPA, proposed by Torbati et al. [119] and built on the MISPEL framework, relaxes the requirement for traveling-subject paired data by employing augmentation strategies on single-site images. Additionally, Tian et al. [112] proposed a bidirectional framework called deep learning-based representation disentanglement (DeRed). This framework consists of four encoders to disentangle anatomical and site-specific representations from paired different sites, and two decoders to bidirectionally synthesize harmonized images. The loss function is composed of four components: (1) Site-consistency loss, which enforce consistent site representations across different subjects from the same site; (2) Subject-consistency loss, which enforces consistent anatomical representations for the same subject scanned at different locations; (3) Self-reconstruction loss, which ensures that the combined representations of the same subject from the same site can reconstruct the original image; (4) Cross-reconstruction loss, which guarantees that cross-site subject pairs can be decoded bidirectionally into images that match their paired counterparts. A key advantage of this model is its flexible multi-site harmonization capability, where new unseen sites can be linked to the target site via intermediate domains without retraining the entire model.
6.3 Available traveling subject datasets
One of the major challenges in image harmonization is how to effectively evaluate its performance. While some studies have employed metrics such as the Fréchet distance, qualitative visual assessment, or feature-level similarity and statistical testing [57]. A more direct and reliable approach is to use traveling-subject datasets, which minimize the bias introduced by inter-site population sampling. However, acquiring such datasets is often costly and limited by the number of available participants. Therefore, to evaluate retrospective image harmonization methods, leveraging existing publicly available traveling-subject datasets is often helpful. Table 2 summarizes the currently available public datasets, covering traveling subjects across different imaging modalities and age groups, and involving major scanner vendors or varying acquisition protocols [110, 125, 116, 109].
Table 2: Available traveling subject dataset
Dataset name Number of subjects (age) Number of Scanners/sites MRI modalities Data Repository
ON-Harmony [125] 20 (18-55y) 6/5 T1w, T2w, SWI, dMRI, rs-fMRI https://openneuro.org/datasets/ds004712
SRPBS [109] 9 (24–32y) 12/8 T1w, rs-fMRI, fieldmap https://bicr-resource.atr.jp/srpbsts
SDSU-TS [51] 9 (22-55 y) 2/2 T1w, T2w, dMRI https://openneuro.org/datasets/ds005664
HAMLET 5 (N/R) 4/3 T1w, dMRI, rs-fMRI https://www.nitrc.org/projects/hamlet
ZJU dMRI [116] 3 (23-26 y) 10/10 T1w, dMRI https://doi.org/10.6084/m9.figshare.8851955.v6
SPINS Human Phantoms [52] 4 (N/R) 6/3 T1w, dMRI, rs-fMRI https://openneuro.org/datasets/ds003011
MUSHAC [110] 14 (21-41 y) 3/ (N/R) dMRI https://www.cardiff.ac.uk/cardiff-university-brain-research-imaging-centre/research/projects/cross-scanner-and-cross-protocol-diffusion-MRI-data-harmonisation
N/R: Not reported
Taking ON-Harmony [125] as an example, 20 healthy volunteers were scanned using five imaging modalities across six scanners from different vendors and models. These modalities included structural imaging (T1-weighted, T2-weighted, and susceptibility-weighted imaging) as well as functional imaging (diffusion MRI and resting-state functional MRI). As shown in Figure 6, a clear observation is that functional modalities exhibit substantially greater inter-scanner variability than structural ones. This discrepancy arises not only from differences in reconstruction and post-processing pipelines across scanners, but also from the fact that both dMRI and fMRI typically rely on fast echo-planar imaging sequences for data acquisition, which are more susceptible to imperfections such as field inhomogeneities and noise.
Figure 6: Representative examples of all modalities for a single participant data across all scanners from ON-Harmony dataset [125].
7 Challenges and Future Directions
Previous research has primarily focused on retrospective harmonization, whereas harmonized acquisition and reconstruction strategies have received comparatively less attention [57, 1, 91]. As an approach that minimizes variability at the source, harmonized acquisition and reconstruction is undoubtedly one of the promising directions for future development. As a representative example, real-time motion tracking and repositioning, which integrates harmonized acquisition with intelligent online reconstruction, has demonstrated prototype-level feasibility in a challenging application like fetal MRI [121, 103]. With the sharing of complete scanning protocols, such strategies can be extended to other applications and have the potential to optimize the entire acquisition workflow in clinical and research settings. However, several limiting factors still need to be addressed before it can be widely adopted. Since this strategy is typically applicable only to newly initiated studies, the size of the available dataset is often constrained. Addressing this limitation requires the sharing of acquisition protocols and source code, as well as active community participation [41, 71]. Another limitation concerns the lack of flexibility in current implementations. For example, vendor support remains limited, online reconstruction frameworks are not yet generalizable, and many open-source platforms still lack flexible user interfaces.
After data acquisition, the choice between image-level and feature-level harmonization represents a fundamental divergence in harmonization strategies. Image-level methods aim to standardize the input data directly, offering broad applicability but potentially at the expense of anatomical fidelity [1]. In contrast, feature-level methods are tailored to specific downstream analyses and are generally safer with respect to anatomical integrity, though they tend to be less generalizable and depend heavily on the chosen features [90]. This strategic decision reflects differing assumptions about the sources of bias and the priorities of the study, such as whether the emphasis is on visual qualitative evaluation or on specific quantitative analyses. This divergence also affects the entire processing pipeline. Image-level harmonization typically occurs early in the workflow, while feature-level harmonization is applied later [57]. As a result, it influences the types of methods used (deep learning is commonly applied at the image level, whereas statistical approaches are more prevalent at the feature level) and the validation required.
The challenge of validating harmonization remains a core bottleneck in the field. The absence of a ground truth and the limitations of current evaluation metrics make the validation process potentially even more difficult than harmonization itself [106, 91]. This hinders the objective comparison of methods and their clinical translation. Developing improved validation strategies and benchmarking frameworks is therefore as critical as developing new harmonization algorithms. Large-scale, multi-modal traveling-subject datasets offer promising opportunities for harmonization validation. However, simply computing similarity metrics between paired images from different sites is not sufficient. Such metrics may overlook subtle but meaningful differences in biological content or downstream analytical relevance, and thus fail to capture the true effectiveness of harmonization. Downstream task-specific harmonization approaches provide a potential solution by aligning harmonization efforts with the objectives of the final application [4, 34, 55, 101]. Nevertheless, these methods require broader exploration and systematic evaluation to ensure generalizability and robustness. Meanwhile, deep learning methods often suffer from a lack of interpretability, the so-called “black box” issue, which makes it difficult to understand how harmonization is achieved and whether the applied transformations are biologically meaningful. Statistical models, while generally more interpretable, often rely on simplified assumptions that may not fully capture data complexity. Hybrid approaches that integrate deep learning with statistical modeling, such as DeepResBat [5] and DeepComBat [58], offer a promising balance between performance and interpretability, and merit further validation and refinement.
Although several large-scale imaging datasets are now publicly available, many targeted neuroimaging investigations remain constrained by small sample sizes and strict data privacy regulations. Traditional harmonization methods often rely on centralized data processing, but the sensitive nature of medical images typically prohibits direct sharing across institutions. Federated learning represents a paradigm shift by enabling each site to train models locally while sharing only model parameters instead of raw data, thereby, in principle, eliminating the need for centralized harmonization [47, 74]. However, directly adapting existing harmonization methods to the federated learning framework is far from straightforward. Many of these approaches require simultaneous access to both source and target domains, or even paired training samples, which is rarely feasible in distributed, multi-site settings [74]. FedHarmony [35] was specifically developed to address these challenges. Like some prior harmonization strategies, it is driven by a downstream predictive task and introduces an auxiliary domain classification objective to explicitly remove site-specific biases. Rather than sharing raw features or images, FedHarmony transmits only statistics, i.e., the mean and standard deviation of the learned features, to update a global knowledge store in a privacy-preserving manner. Local models are trained using adversarial domain adaptation and are then aggregated using a site-balanced strategy. In doing so, FedHarmony effectively removes scanner-specific effects while maintaining consistent task performance across sites, demonstrating that harmonization within federated learning is not only feasible but also a promising and privacy-conscious direction for multi-site neuroimaging research.
An alternative strategy for addressing data privacy concerns and mitigating data bias introduced by undesired factors is to explicitly simulate such “biases” during training, as represented by synthetic data-driven deep learning strategies [134, 133, 124]. Synthetic data are typically generated in large volumes using predefined anatomical priors or physical models, with randomized variations introduced in a controlled manner. This randomization substantially expands the diversity of the data distribution, thereby enhancing the generalizability of the trained models to unseen scenarios [45]. Crucially, because synthetic data are inherently devoid of site-specific biases, models trained on such data may promote harmonization when applied to multi-site datasets. This phenomenon was preliminarily demonstrated in the study by Sun et al. [108], where a foundation model trained on synthetic data was developed to enhance MR images by simultaneously correcting for artifacts, suppressing noise, and improving resolution. When applied to real-world data acquired from different scanners and sites, the output images exhibited more consistent contrast and histogram distributions, indicating that the foundation model had effectively learned to project heterogeneous inputs into a shared representational space. In addition, synthetic data can facilitate harmonization through unified domain translation. For example, SynthSR [60] harmonizes images acquired under varying field strengths, modalities, and acquisition protocols by converting them into a standardized T1-weighted structural representation. SynthSeg [14] further extends this approach by harmonizing multi-modal inputs for disease-specific segmentation and analyses, thereby enabling the reuse of legacy datasets.
8 Conclusion
In this survey, we reviewed the current advanced methods for medical image harmonization. Unlike previous review papers, we firstly include the emerging and promising direction of harmonized image acquisition, highlighting accessible open-source tools that support its implementation. For post-acquisition harmonization, we focused on deep learning-based approaches, categorizing representative methods into two major groups: image-level and feature-level strategies. We also summarized the available traveling-subject datasets that support analyses and validations. Furthermore, we discussed several forward-looking directions including intelligent acquisition repositioning, federated learning frameworks, foundation models, and the use of synthetic data. This survey provides a comprehensive reference for researchers, and supports the continued development of harmonization techniques in multi-site, multi-modal medical imaging.
Acknowledgments This research was supported in part by the National Institute of Health (NIH) under award numbers R01EB031849, R01EB032366, and R01HD109395. The content of this publication is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Data availability No data was used for the research described in the article.