### v2

# Harmonization in Magnetic Resonance Imaging: A Survey of Acquisition, Image-level, and Feature-level Methods

Abstract
Magnetic resonance imaging (MRI) has greatly advanced neuroscience research and clinical diagnostics. However, imaging data collected across different scanners, acquisition protocols, or imaging sites often exhibit substantial heterogeneity, known as “batch effects” or “site effects.” These non-biological sources of variability can obscure true biological signals, reduce reproducibility and statistical power, and severely impair the generalizability of learning-based models across datasets. Image harmonization is grounded in the central hypothesis that site-related biases can be eliminated or mitigated while preserving meaningful biological information, thereby improving data comparability and consistency. This review provides a comprehensive overview of key concepts, methodological advances, publicly available datasets, and evaluation metrics in the field of MRI harmonization. We systematically cover the full imaging pipeline and categorize harmonization approaches into prospective acquisition and reconstruction, retrospective image-level and feature-level methods, and traveling-subject-based techniques. By synthesizing existing methods and evidence, we revisit the central hypothesis of image harmonization and show that, although site invariance can be achieved with current techniques, further evaluation is required to verify the preservation of biological information. To this end, we summarize the remaining challenges and highlight key directions for future research, including the need for standardized validation benchmarks, improved evaluation strategies, and tighter integration of harmonization methods across the imaging pipeline.
keywords:
Magnetic Resonance Imaging , Image Harmonization , Deep Learning
1 Introduction
Magnetic resonance imaging (MRI) has had a profound impact on medicine, with widespread applications in medical and neuroscience research, computer-aided diagnosis, longitudinal monitoring, and image-guided interventions. To advance scientific discovery and bridge the gap between research and clinical practice, the collection and sharing of large-scale imaging datasets across sites has become increasingly essential [128, 125, 88, 115, 110, 21]. Multi-center studies that aggregate large and diverse samples not only enhance statistical power, particularly important for investigating rare or low-prevalence diseases, but also provide broader coverage of key biological variables such as age, sex, race, geographic location, socioeconomic status, and disease subtypes. The increased sample size and heterogeneity also improve the ability of studies to detect subtle yet meaningful effects in high-dimensional spaces of variables and confounders [89, 15].
In the era of deep learning, the proper collection and analysis of large-scale medical imaging data has become even more critical [142, 81, 90]. Although machine or deep learning models have shown great potential in addressing scientific challenges in medicine, their applicability in clinical practice remains limited. Many of these models are built on the assumption that the training and testing datasets come from the same distribution, and their performance can degrade substantially when this assumption does not hold. In other words, when applied to images acquired using protocols that are different from those used during training, these models often exhibit poor reproducibility and limited generalizability [50]. As a result, their effectiveness may decline significantly when used across different hospitals, scanners, or patient populations. This decline in performance is largely caused by non-biological technical variability, often referred to as scanner effects, device effects, or batch effects, which can stem from differences in scanner/device hardware and software by different manufacturers, sequences, acquisition protocols, image reconstruction pipelines and techniques, and other sources [86, 27, 64, 53].
A common conventional approach to avoid the challenges of directly using and comparing heterogeneous imaging data is meta-analysis, where each site performs its own analysis independently, and the results are later combined [3, 74, 36]. However, meta-analyses typically include only group-level statistical and clinical information, making it difficult to perform detailed modeling or adjustments at the individual level. Moreover, when participant distributions are imbalanced, site-specific statistics may introduce systematic biases. In studies with small imaging sample sizes, fluctuations in parameter estimation during z-score conversion may further compromise the stability of statistical inference. In contrast, mega-analysis involves the joint analysis of all raw imaging data on a unified platform [89, 85, 29, 62, 58]. However, this strategy places more stringent demands on data harmonization, as combining datasets from different centers may introduce additional non-biological variability, particularly due to differences in imaging protocols. Therefore, effective harmonization is a critical prerequisite for the success of mega-analyses.
Comprehensive MRI harmonization involves three key components: harmonized acquisition, image-level processing, and feature-level analysis. Harmonized acquisition refers to the prospective standardization of the use of scanner hardware, pulse sequences, and protocol parameters during data collection, aiming to reduce variability or heterogeneity of the acquired data [77]. Image-level harmonization involves retrospective adjustments to the acquired images, such as intensity normalization, statistical correction, or deep learning methods to standardize image contrast and signal distribution [62, 1, 37]. Feature-level harmonization focuses on quantitative metrics, texture features, and anatomical representations extracted from images, ensuring their comparability across different sites to support reliable data integration and cross-site analysis [43, 141, 42]. The primary goal of harmonization is not to recover some absolute “ground truth,” but rather to enhance the reliability of comparisons at both individual and group levels. In other words, the goal of harmonization is to reduce non-biological technical variation while preserving meaningful biological differences. Rather than requiring complete elimination of systematic biases, many harmonization approaches aim to make such biases consistent across all datasets, thereby minimizing their impact on downstream analyses.
This review focuses on multi-site harmonization methods for MRI data, although the underlying principles are broadly applicable to other medical imaging modalities. Several review papers have been published on this topic, covering modalities such as positron emission tomography, computed tomography, and microscopic pathology [62, 1, 97, 131, 11, 94]. However, the primary emphasis remains on MRI, owing to its inherent characteristics, such as multiple field strengths, diverse imaging modalities, and a wide range of quantitative parameters, which result in pronounced inter-site heterogeneity. Nevertheless, existing MRI harmonization reviews have largely focused on retrospective approaches applied to structural and diffusion MRI [62, 1, 97], while prospective strategies at the acquisition stage have received limited attention. In this review, we place particular emphasis on recent advances in vendor-agnostic pulse sequences and harmonized image reconstruction techniques. For retrospective harmonization, we highlight emerging deep learning-based methods that have gained traction in recent years. This is the first comprehensive review to cover the full spectrum of harmonization efforts across the MRI pipeline, from acquisition to analysis, spanning prospective acquisition harmonization and retrospective image- and feature-level methods (Figure 1). Rather than providing an exhaustive survey, we focus on representative methods that capture the core ideas of each category.
Figure 1: Overview of harmonization strategies across the entire MRI pipeline, covering image acquisition, reconstruction, preprocessing, feature extraction and analysis, and including representative methods discussed in this review. dMRI, diffusion-weighted MR imaging; T1WI, T1-weighted imaging; T2WI, T2-weighted imaging; RBV, regional brain volume; CT, cortical thickness; FA, fractional anisotropy; TS, Traveling subject.
2 Background
Table 1: Sources of measurement and sampling bias in multi-site MRI data
Measurement Bias (non-biological) Sampling Bias (biological)
Scanner properties (hardware): - Scanner manufacturer - Field strength - Gradient systems - Radiofrequency coils - Hardware imperfections - Field inhomogeneity ($\Delta B_{0}$, $B_{1}^{+}$, $B_{1}^{-}$) Demographic and population: - Age - Sex - Population background (e.g., ethnicity, genetics)
Pulse sequence design (software): - Tissue contrast sensitivity - Motion and flow sensitivity - Vendor-specific implementations General health status: - Anthropometric/physiological measures    (e.g., BMI, blood pressure) - Co-occurring disorders - Lifestyle and environmental factors
Acquisition parameters (software): - Spatial resolution: FOV, matrix size, slice thickness - Signal-to-noise ratio (SNR): flip angle,    receiver bandwidth, number of excitations - Contrast parameters: TE, TR, TI,    preparation modules, b-value, gradient directions Clinical characteristics: - Disease presence, subtype and severity - Symptom heterogeneity - Medication and Treatment history
Image post-processing (software): - Image normalization and filtering - Image reconstruction and multi-coil combination - Distortion and artifact correction Study protocol: - Subject inclusion criteria - Diagnostic criteria and instruments - Case-control matching strategies
Abbreviations: BMI, body mass index; FOV, field of view; TE, echo time; TR, repetition time; TI, inversion time.
Site differences are more pronounced in diffusion and functional MRI due to echo-planar imaging (EPI)-induced effects, including odd-even phase differences, geometric distortions, and signal loss, as well as variations in acquisition geometry, brain coverage, task implementation, and temporal duration.
2.1 Sources of data heterogeneity
The collection and analysis of MRI data involve a complex and variable set of procedures, including subject recruitment/selection, imaging hardware, protocol design, and the choice of data analysis models. Variations at any of these stages (across imaging sites, subjects, or longitudinal scans) can introduce differences into the final results [39, 8, 101]. Therefore, to achieve effective data harmonization, i.e., the removal of non-biological variability while preserving biologically meaningful differences across participants and sites, it is essential to understand the sources of such variability. The major sources contributing to measurement and sampling bias are summarized in Table 1.
2.2 Prospective and retrospective approaches
Prospective harmonization refers to strategies that are deliberately planned prior to data acquisition with the goal of minimizing anticipated sources of variability at the source [29, 62, 77]. A common approach involves standardizing scanner models and acquisition protocols in advance to reduce differences introduced during image acquisition. Building on this foundation, this review highlights recent advances such as vendor-agnostic pulse sequences and harmonized image reconstruction methods, which aim to overcome traditional barriers to acquisition consistency through open-source and easily implementable solutions. These innovations further enhance the effectiveness of prospective harmonization. In addition, the use of traveling subjects represents another important prospective strategy, whereby the same individuals are scanned across multiple sites to obtain matched datasets. This design provides a valuable reference for quantifying and correcting systematic inter-site differences during analysis, thereby supporting the effective removal of non-biological variability.
Retrospective harmonization refers to the process of applying harmonization techniques after data acquisition to existing, heterogeneous multi-site datasets, with the aim of eliminating or significantly reducing variability caused by non-biological factors [62, 97, 91, 71]. These methods currently dominate the field, largely due to the accessibility of large-scale public multi-site datasets and their advantages in terms of cost-effectiveness and flexibility compared to prospective approaches. Current retrospective strategies include traditional image processing, statistical modeling, and deep learning-based methods. According to the object of harmonization, they can be divided into image-level and feature-level strategies, as summarized in Section 2.3. Their effectiveness requires validation from multiple complementary perspectives (Section 7), and a unified evaluation standard has yet to be established.
2.3 Image-level and feature-level approaches
Image-level harmonization aims to directly modify the voxel intensities of MR images and is typically framed as an image-to-image translation task. The goal is to make images acquired from different sources be similar in terms of characteristics such as contrast, sharpness, and signal-to-noise ratio (SNR), as if they were acquired under comparable settings. In principle, harmonized images can then be used for a variety of downstream tasks such as segmentation and feature extraction, and may also be visually inspected by radiologists. However, image-level harmonization carries the risk of altering underlying anatomical structures or introducing artifacts, especially when using complex generative models based on deep learning. Therefore, it is crucial to properly validate techniques by validating the biological fidelity of the harmonized images [62, 37, 131, 94].
Feature-level harmonization methods operate on quantitative features or image-derived measurements extracted from raw MRI data, such as regional brain volumes, cortical thickness, diffusion tensor imaging (DTI) metrics, functional connectivity matrices, or radiomic features. This makes them well-suited for incorporating statistical models to remove site effects. One key advantage of feature-level approaches is their ability to directly include biological covariates in the modeling process. Operating on derived features rather than raw images reduces the risk of altering image appearance and improves computational efficiency, but it also limits data reusability and makes harmonization performance dependent on the feature extraction procedure [62, 43, 5, 63].
3 Harmonized Data Acquisition
An MRI system consists of two main components: hardware and software (Figure 1). Although exact matching of both components would ideally enable optimal multi-site harmonization, this is rarely achievable in large-scale studies, particularly in international collaboration. Compared with hardware, harmonizing software offers greater flexibility and can be categorized into vendor-agnostic acquisition and harmonized image reconstruction approaches.
3.1 Vendor-agnostic pulse sequence
Since pulse sequences serve as the core of MR image formation, their harmonization offers a principled approach to addressing site effects at the source. However, due to differences in the underlying implementation of pulse sequences from different vendors, signal discrepancies may still arise even when identical acquisition parameters listed in Table 1 (e.g., TE, TR, FOV, matrix size) are used [70]. Such variations originate from factors including, but not limited to, differences in fat suppression modules, dephasing strategies (e.g., spoiler and crusher gradients), RF pulse shapes and profiles, most of which are not accessible or adjustable through the user interface [77, 45]. To address this challenge and enhance consistency at sequence level, several vendor-agnostic or open-source pulse sequence platforms have been developed over the past decade, including Pulseq [77], gammaSTAR [76] and RTHawk [105]. For example, Pulseq enables modular pulse sequence programming in MATLAB and Python, allowing extensive and detailed control over RF pulses, gradient waveforms, and inter-module interval. The resulting sequence is compiled into a standardized .seq file, which can then be interpreted and executed by vendor-specific backends for MRI scanning (Figure 2). Additionally, Pulseq can be integrated with various MRI simulation and graphical sequence design tools, further alleviating the steep learning curve of pulse sequence development [7].
Figure 2: Harmonized acquisition and reconstruction workflow proposed in [45]. The pulse sequence was implemented using Pulseq to ensure identical configurations across scanners and vendors. All post-processing steps including image reconstruction and quantitative parameter fitting were performed offline using a consistent pipeline.
Liu et al. [82] systematically evaluated a single-shell diffusion MRI sequence implemented using Pulseq across two scanners from different vendors, using standard error as a metric of repeatability. For mean diffusivity in phantoms, the Pulseq sequence demonstrated 2.5-fold superior inter-scanner reproducibility compared to vendor-provided sequences. In human brain imaging, a Pulseq sequence reduced inter-scanner standard error in fractional anisotropy by 35–50% across various brain regions. In addition to diffusion MRI, vendor-agnostic pulse sequence tools have also been validated in quantitative MRI (qMRI) applications, including chemical exchange saturation transfer [59], brain T1 and T2 mapping [45, 73], and myocardial T1 mapping [49, 48]. In the work by Karakuzu et al. [70], combining RTHawk-based acquisition harmonization with a unified parameter quantification workflow led to statistically significant reductions in inter-vendor variability for T1, magnetization transfer ratio, and magnetization transfer saturation index measurements. Although current preliminary results are encouraging, it remains unclear to what extent vendor-agnostic or open-source pulse sequence tools can mitigate site effects, as they have yet to be widely implemented or validated at scale.
3.2 Harmonized data reconstruction
The raw MRI signal acquired from the scanner is one-dimensional complex-valued data. To generate the final image, this signal should be filled into a predefined k-space trajectory and then transformed via Fourier transformation, which is known as image reconstruction. Differences in reconstruction pipelines can introduce non-negligible variability, contributing to a lack of harmonization across sites or vendors. These differences may arise from multiple factors, including pre- and post-reconstruction distortion and phase correction, k-space gridding, partial Fourier reconstruction, multi-coil parallel reconstruction, and coil combination strategies [54]. While vendors provide access and control over some of the options, filters, and parameters through user interfaces during acquisitions, a full control over the entire raw data processing pipeline often requires additional programming within the vendor software environment, which may not always be available or if available, may not be straightforward.
Offline open-source reconstruction toolboxes offer alternative, promising opportunities for standardizing the image reconstruction process. Representative examples include the Berkeley Advanced Reconstruction Toolbox (BART) [124] and the Michigan Image Reconstruction Toolbox (MIRT) [41]. For instance, BART not only implements conventional parallel imaging algorithms but also provides general-purpose solutions for non-Cartesian, model-based, and deep learning-based reconstruction [18]. Its cross-platform, open-source, and multi-language support (Linux terminal, MATLAB, and Python) make it accessible and easy to integrate into diverse research workflows. Prior to reconstruction, inconsistencies in raw data formats across vendors pose a significant challenge. The ISMRMRD (International Society for Magnetic Resonance in Medicine Raw Data) [67] framework addresses this issue by providing a standardized format that harmonizes vendor-specific raw data and headers, thereby facilitating consistent and reproducible reconstruction pipelines.
Despite their ease of use, offline open-source reconstruction toolboxes have inherent limitations that restrict their clinical scalability. These include the lack of real-time quality control and the requirement to store large raw datasets. To address these challenges, online reconstruction frameworks, e.g., Gadgetron [55] and FIRE (Framework for Image Reconstruction Environments) [32, 9], have been proposed. Gadgetron adopts a modular, streaming-based architecture and incorporates a wide range of extensible toolboxes, enabling real-time reconstruction through GPU or multithreaded CPU acceleration. It also supports advanced features such as automated motion tracking and scan planning, thereby minimizing heterogeneity arising from operator-dependent variability across sites. A notable example is the HERON framework for online acquisition and reconstruction of fetal diffusion-weighted MR imaging (dMRI) based on real-time motion quantification [126]. This approach leverages image-based real-time motion estimation to dynamically adjust fetal dMRI acquisition, thereby mitigating the impact of unpredictable fetal motion. In this context, Gadgetron is used to enable real-time landmark detection through deep learning and motion analysis, providing dynamic feedback for prospective adjustment of the acquisition. Similar strategies have also been extended to fetal functional MRI (fMRI), qMRI, and automated cardiac scan planning [108, 109, 127, 20, 17]. Online reconstruction in these challenging applications enables motion-informed data re-acquisition, further promoting data consistency across scans and sites.
3.3 Modality applicability and limitations
Harmonized data acquisition and reconstruction are, in principle, applicable to nearly all MRI modalities, as they are inherently prospective approaches. Existing efforts have primarily focused on qMRI, dMRI, and fMRI, particularly those involving advanced pulse sequences [28, 117, 104, 84]. Vendor-neutral acquisition provides a unique avenue for disentangling the sources of variability in quantitative and functional MRI, enabling systematic investigation of whether observed differences arise from physiological factors or technical confounds. In addition, open-source frameworks facilitate rapid validation and dissemination of new methods, improve methodological transparency, and thereby promote reproducible cross-site studies. Taking Pulseq as an example, the community-driven Harmonized MRI [^1] https://harmonizedmri.github.io/projects/ initiative has already collected dozens of open-source projects, and recent studies increasingly adopt the practice of publicly releasing their acquisition sequences and reconstruction pipelines alongside publication [28, 117].
Despite these advantages, the limitations of harmonized data acquisition and reconstruction should not be overlooked. At a conceptual level, such approaches provide a principled means of controlling variability at its source, yet are not directly applicable to existing large-scale and heterogeneous MRI datasets. In practice, current vendor-neutral pulse sequences often provide limited parameter flexibility and interactivity: once a sequence file is generated, most acquisition parameters are difficult to adjust in response to practical scanning conditions. Moreover, both the development and practical deployment of open-source sequences and real-time reconstruction remain dependent on vendor-level support, and their integration into standard clinical workflows requires substantial effort and faces regulatory barriers. As a result, these constraints hinder the adoption of harmonized acquisition and reconstruction as scalable clinical or multi-center solutions, confining their use largely to prospective, small-scale studies
4 Image-level Retrospective Harmonization
4.1 Traditional image-level approaches
Early traditional image-level harmonization methods primarily rely on various intensity normalization techniques [96, 107]. Although some of these methods are not explicitly designed to remove site effects, they are commonly used as preprocessing steps or baseline approaches due to their simplicity and low computational cost. Such methods typically apply global transformations to the entire image (e.g., z-score normalization), adjust image intensity statistics (e.g., histogram matching), or utilize reference intensities from specific tissue types to align global or local intensity distributions, thereby improving comparability across scans acquired from different sites.
A widely used class of methods is based on histogram matching [96, 106], which aims to align the intensity histogram or cumulative distribution function (CDF) of a source image with that of a target image or a predefined reference distribution. Although these methods are conceptually simple and computationally efficient, they are often sensitive to outliers (e.g., hyperintense lesions) and may fail to preserve biologically meaningful variations at the individual level. Another class of methods relies on reference-based normalization, such as White Stripe proposed by Shinohara et al. [107], which rescales the image intensities using a reference region composed of normal-appearing white matter (NAWM). Building upon this approach, RAVEL (Removal of Artificial Voxel Effect by Linear regression) [44] further addresses residual non-biological variability that may persist after White Stripe normalization, which will be introduced in Section 5.1. However, White Stripe relies on the assumption that NAWM serves as stable reference, which may not hold in populations with white matter pathology.
For diffusion MRI data, harmonization can be performed directly on the diffusion-weighted (DW) signal prior to feature extraction. Cross-site DW data are often affected by differences in b-values, the number of diffusion gradient directions, angular sampling density, and global signal scaling induced by hardware or reconstruction/phase correction differences [97, 43]. These discrepancies cannot be adequately modeled as a simple intensity offset. A representative approach for dMRI is RISH (Rotationally Invariant Spherical Harmonics) [91, 71], which decomposes voxel-wise DW signals into spherical harmonics for harmonization. In practice, RISH extracts spatially varying voxel-wise scaling factors by aligning RISH features across matched control groups and applies these factors to individual DW data. Due to its rotational invariance, RISH are robust to differences in gradient orientations and can serve as a preprocessing step compatible with a wide range of downstream analysis pipelines. De Luca et al. [35] further extended the RISH framework by introducing a covariate-driven general linear model (RISH-GLM), allowing multivariate modeling of site effects and cross-site harmonization without the need for matched reference data. Another representative approach is the Method of Moments (MoM) proposed by Huynh et al. [65]. MoM derives voxel-wise scaling parameters by matching the spherical mean and variance of DW signals across sites and applies them to achieve signal-level harmonization. Compared with RISH, this method can be flexibly applied to datasets acquired with different numbers of gradient directions.
4.2 Learning-based image-level approaches
Learning-based image-level harmonization methods are predominantly driven by deep convolutional neural networks and can be broadly categorized into four groups: adversarial learning and style transfer, anatomy-contrast disentanglement, multi-contrast prior learning, and source-free distribution modeling. A general trend across these approaches is the shift from fixed, site-specific harmonization toward adaptive multi-site solutions, and from methods requiring simultaneous access to data from multiple sites to those leveraging only single-site data.
4.2.1 Adversarial learning and style transfer
For learning-based approaches, Image-level harmonization aligns with the fundamental characteristics of conventional image style transfer. Consequently, generative adversarial networks (GANs) and their variants, as leading techniques in style transfer tasks, were among the first methods employed to address the challenge of unpaired image harmonization [144] (Figure 3a), with Cycle-consistent GAN (CycleGAN) being the most widely used approach [92, 118, 25, 99]. The basic strategy involves treating unpaired data from two different sites as the source and target domains, respectively, and applying cycle-consistency loss to transfer style information while preserving biological structures. Building on this foundation, improvements to the generator/discriminator architecture, modifications to loss functions, or the incorporation of attention mechanisms have been explored to enhance training stability and harmonization performance.
Figure 3: Four representative categories of image-level deep learning-based harmonization methods: (a) adversarial learning and style transfer, (b) anatomy-contrast disentanglement, (c) multi-contrast prior learning and (d) source-free distribution modeling. A: anatomy; S: style; C: content.
A major limitation of CycleGAN is that it requires separate training for each pair of sites and fails to leverage complementary information across multiple sites. To address this, several many-to-one GAN-based harmonization strategies have been proposed [103]. For example, the IGUANe framework proposed by Roca et al. [103] employs a single universal generator capable of translating images from multiple source sites into the style of a reference site. In practice, this universal generator is adversarially trained against multiple site-specific discriminators, while multiple backward generators are trained similarly to enforce cycle consistency. In a harmonization task involving 11 sites/scanners, the IGUANe architecture required training 11 generators and 20 discriminators, resulting in substantial computational overhead. StyleGAN [72] and StarGAN [30, 31] further extend CycleGAN by introducing explicit site or style encodings, which significantly reduce the number of required encoders and discriminators. For instance, Bashyam et al. [10] achieved harmonization across six sites using only four models based on the StarGAN framework, and demonstrated improved accuracy in brain age prediction in downstream tasks.
4.2.2 Anatomy-contrast disentanglement
GAN-based image style transfer offers a straightforward solution for harmonization, but it lacks explicit separation between anatomical structures and contrast-related features. In contrast, variational autoencoder (VAE), which encode images into low-dimensional latent representations and decode them back to images, can be used to explicitly disentangle anatomy and contrast (Figure 3b).
In the MURD method proposed by Liu et al. [83], independent style encoders, content encoders, and image generators are introduced. Once the latent features representing anatomy and style are explicitly disentangled, image harmonization can be achieved by recombining the anatomical representation of the source domain with the style of the target domain. This strategy not only accommodates resolution differences across sites, but also enables continuous harmonization of imaging styles between sites. In contrast, ImUnity, proposed by Cackowski et al. [22], adopts contrastive learning to simplify training. Anatomical and style disentanglement is achieved by leveraging anatomical similarity across different slices from same subject together with gamma-based intensity transformations. Furthermore, DLEST, proposed by Wu et al. [135], performs harmonization directly in latent space. Site-invariant image generation is first learned from single-site data to obtain encoder–decoder pairs for latent extraction and image reconstruction. Latent representations from different sites are then aligned using an energy-based model. Harmonized images are finally generated by decoding the transformed latent representations.
4.2.3 Multi-contrast prior learning
Clinical acquisitions and large-scale public MRI datasets often include multi-contrast images, such as T1, T2, PD (proton density) and FLAIR (fluid attenuated inversion recovery), to capture complementary tissue properties. It is commonly assumed that these contrasts, when acquired from the same subject, share a consistent anatomical structure, thereby providing naturally paired anatomy-contrast information for disentanglement (Figure 3c). This mitigates concerns regarding pathology or demographic differences between the groups and reduces, or even eliminates, the need for simultaneous access to both source and target datasets.
Dewey et al. [38] used co-registered T1- and T2-weighted images from the same subject as inputs to an encoder to disentangle the latent space. During inference, images of arbitrary contrast from a new site can be encoded, combined with a fixed contrast vector from a reference site, and decoded to produce harmonized images. The CALAMITI method [146], proposed by the same group, introduces cross-contrast synthesis and adversarial learning across different sites. These enhancements enable CALAMITI to achieve more robust disentanglement and to learn a globally consistent anatomical representation across sites. Furthermore, the HACA3 method proposed by Zuo et al. [147] incorporates more than two MRI contrasts (T1, T2, PD and FLAIR) and supports flexible combinations of arbitrary modalities, making it well-suited for scenarios where multi-contrast data may be incomplete or heterogeneous.
In addition to the structural similarity, multi-contrast MR images are linked through inherent physical properties. Specifically, different contrast-weighted images acquired via varying pulse sequences and acquisition parameters ultimately reflect the same underlying tissue properties, i.e. the quantitative T1 and T2, and PD. Based on this, Qiu et al. [100] and Borges et al. [19] proposed physics-driven harmonization frameworks that translate multi-contrast MRI into quantitative maps that are invariant to modality and acquisition protocol. In these approaches, contrast-specific forward physical models are embedded into the loss function to enable self-supervised learning. Additionally, PhyCHarm [78] further leverages scanner-specific acquisition parameters to synthesize native MR images from the quantitative maps for harmonization. However, the final step still requires supervision from paired training data.
4.2.4 Source-free distribution modeling
To eliminate the dependency on multi-contrast and multi-site data while ensuring generalizability to unseen domains, a new generative model, normalizing flow, was introduced to directly model the source distribution. This enables source-free harmonization without the need for traveling subjects, multi-contrast data, or task-specific supervision, as the model learns to map a complex probability distribution to a simple one through a series of invertible and differentiable transformations.
Jeong et al. and Beizaee et al. independently and almost simultaneously introduced normalizing flows into the field of image harmonization, proposing the BlindHarmony [68] and Harmonizing flows [13] frameworks, respectively. Taking Harmonizing flows as a representative example, the method adopts a three-step harmonization strategy consisting of source domain modeling, harmonizer pre-training, and test-time domain adaptation. Specifically, a normalizing flow network, composed of stacked affine coupling layers, was first trained to capture the distribution of source domain images. This model provides an invertible and differentiable transformation that maps the complex source distribution to a standard Gaussian. Subsequently, a lightweight UNet was employed as the harmonizer and pre-trained to reconstruct original source images from their randomly augmented counterparts, thereby learning to compensate for appearance variations introduced by contrast and intensity shifts. Finally, during inference, the harmonizer was fine-tuned using target domain images under the supervision of the frozen flow model, ensuring that the harmonized outputs align with the source domain distribution (Figure 3d). This framework achieves unsupervised, source-free and task-agnostic harmonization, and demonstrates generalizability to unseen domains, with robust performance across multiple tasks, including brain MRI segmentation and neonatal age estimation.
4.3 Modality applicability and limitations
Image-based harmonization methods have been predominantly developed for conventional structural MRI contrasts, such as T1, T2, PD, and FLAIR images. Although a limited number of studies have claimed potential generalizability to other MRI modalities, substantive evidence supporting such claims remains scarce. This lack of validation underscores the limited generalizability and cross-modality transferability of image-based harmonization approaches. In the context of dMRI, harmonization efforts are dominated by RISH-based methods and their extensions, whereas learning-based approaches largely rely on traveling-subject data for supervision (Section 6.2). In contrast, to date, no dedicated image-based harmonization methods have been established for fMRI or qMRI.
Beyond their limited generalizability, image-based harmonization methods are also prone to over-correction and the removal of biologically meaningful variability. This risk primarily arises from the lack of explicit covariate constraints and the failure to account for the confounding between site effects and biological variables. While learning-based approaches have substantially advanced harmonization performance, these gains come at the cost of reduced interpretability, increased dependence on large-scale training data and computational resources. Such methods may also introduce hallucination effects that are inherently challenging to detect and avoid. This limitation is exacerbated by the fact that existing validation strategies largely rely on qualitative visual comparisons of samples outside the training data [13]. Furthermore, image-based methods are inherently sensitive to preprocessing pipelines and registration accuracy, and their harmonization outcomes are often strongly dependent on the choice of reference site.
5 Feature-level Retrospective Harmonization
5.1 Statistical approaches
Feature-level methods based on statistical modeling typically assume that extracted features (e.g., brain volumes, cortical thickness, DTI metrics) can be decomposed into biological effects, site or batch effects, and random noise. By fitting a statistical model (most commonly a linear model), the site effect can be estimated and subsequently removed or adjusted, yielding harmonized data. A comprehensive review of such methods is available in Hu et al.[62] Here, we briefly introduce several representative approaches, with a particular emphasis on ComBat [43, 69], which serves as a foundation for subsequent learning-based methods.
The ComBat model was originally developed for batch effect correction in gene expression data [69], and was first introduced to DTI data [43] In this model, the observed feature $y_{ijv}$ at voxel $v$ for subject $j$ from site $i$ is modeled as a linear combination of multiple factors. These factors include the global mean $\alpha_{v}$, biological covariates (e.g., age and sex), and site effects (additive and multiplicative effects, $\gamma_{iv}$ and $\delta_{iv}$). The full model can therefore be expressed as:
$y_{ijv}=a_{v}+\mathbf{X}_{ij}\beta_{v}+\gamma_{iv}+\delta_{iv}\varepsilon_{ijv}$ (1)
where $\mathbf{X}$ is the design matrix for the covariates, $\beta$ represents the corresponding regression coefficients, and $\varepsilon$ is the error term, assumed to have zero mean and variance $\sigma^{2}$. After estimating the coefficients using the empirical Bayes method, the ComBat-harmonized value can be expressed as:
$y_{ijv}^{\mathrm{ComBat}}=\frac{y_{ijv}-\hat{a}_{v}-\mathbf{X}_{ij}\hat{\beta}_{v}-\hat{\gamma}_{iv}}{\hat{\delta}_{iv}}+\hat{a}_{v}+\mathbf{X}_{ij}\hat{\beta}_{v}$ (2)
Based on this, ComBat has been shown to be effective across a variety of imaging features, including not only DTI-derived metrics but also cortical thickness, functional connectivity, spectroscopy and radiomics [141, 42, 101, 132, 2, 14, 75]. It is worth noting that, although ComBat can also be applied after normalizing images to a standard space (image-level harmonization), this is often a suboptimal choice and is therefore typically used only as a comparative baseline [13, 60].
Based on the standard ComBat, numerous extensions have been proposed, including the use of alternative parameter estimation strategies and applications to more complex study designs [98, 61, 102, 121, 23, 33, 145, 137]. For example, ComBat-GAM introduces generalized additive models (GAMs) to model nonlinear covariate effects on feature means [98]. ComBatLS further generalizes this approach using GAMLSS (generalized additive models for location, scale, and shape) to account for covariate-dependent variability in both mean and variance [47]. However, one major limitation of ComBat-based methods is their reliance on sufficient within-scanner samples to estimate site-specific effects. This constraint reduces its generalizability to unseen data [5].
Unlike ComBat, which explicitly models additive and multiplicative effects, some other strategies model biological factors or site effects using basis representation or latent factors [44, 40, 26, 79, 143]. Taking CovBat [26] as an example, it extends ComBat by explicitly accounting for site effects in covariance. CovBat performs principal component analysis on the ComBat-adjusted residuals and conducts harmonization in the principal component space, yielding CovBat residuals, $\hat{\varepsilon}_{ijv}^{\mathrm{CovBat}}$. Consequently, Eq. (2) is reformulated as:
$y_{ijv}^{\mathrm{CovBat}}=\hat{\varepsilon}_{ijv}^{\mathrm{CovBat}}+\hat{a}_{v}+\mathbf{X}_{ij}\hat{\beta}_{v}$ (3)
Similarly, Zhang et al. proposed FELIEF (REmoval of Latent Inter-scanner Effects through Factorization) [143], which addresses more complex site effects by explicitly modeling scanner-related latent multivariate structures. The method applies matrix factorization to standardized residuals and imposes low-rank constraints on the latent components via nuclear norm regularization, thereby effectively identifying and removing scanner-related technical variability. Another representative example is RAVEL [44], which selects a control voxel that is highly sensitive to variations in reconstruction algorithms, acquisition protocols, and scanner configurations, typically from the cerebrospinal fluid, to serve as a proxy for non-biological effects. Then, RAVEL performs singular value decomposition on the control voxels to extract latent factors representing technical variation, and then applies linear regression across all voxels to estimate and remove these effects. Finally, the resulting residuals are treated as the RAVEL-corrected intensities.
5.2 Learning-based feature-level approaches
Learning-based feature-level approaches are typically extensions of statistical harmonization strategies, particularly those derived from ComBat. A representative example is Neuroharmony [46, 6], which is based on the assumption that the intrinsic image characteristics of a single image can aid in data harmonization. This approach addresses the limitation of traditional ComBat, which cannot generalize to unseen sites. Specifically, Neuroharmony first applies ComBat to existing multi-site data to obtain corrected features for each image. Then, using the MRIQC tool, a range of image quality metrics (IQMs) are extracted from each image, including SNR, contrast, blurriness, motion artifacts, and background uniformity. Based on these metrics, Neuroharmony employs a random forest model to learn the mapping between the 64 IQMs and the ComBat-derived corrected features. Once trained, the model no longer relies on population-level statistical features but instead performs harmonization using only the image’s IQMs and biological covariates.
Another line of learning-based feature-level approaches leverages conditional variational autoencoder (cVAE) [93] to address nonlinear site-related variations and support multivariate modeling. In the cVAE framework, an encoder first processes feature vectors to generate latent representations. These representations are then concatenated with site information (i.e., a one-hot vector) or biological covariates and fed into a decoder to reconstruct the original feature vectors. To accommodate 1D input, both the encoder and decoder are typically implemented as fully-connected neural networks. To encourage site-invariant latent representations, mutual information between the latent features and the site encodings is minimized during training. Then, in the harmonization phase, modifying the input site encoding allows for flexible translation of input features to any target site.
Several extensions of the cVAE framework have been developed to enhance harmonization performance. For instance, the goal-specific cVAE (gcVAE) [4] proposed by An et al. incorporates a pretrained classifier into the standard cVAE architecture. This allows the original cVAE to implicitly preserve biologically meaningful representations by leveraging supervision from downstream classification tasks during training. Another variant, DeepComBat by Hu et al. [63], integrates cVAE with the classical ComBat method. It first applies ComBat to the latent mean vectors produced by the cVAE encoder, followed by decoding the harmonized latent representations to reconstruct the original features. A second ComBat step is then applied to the residuals (i.e., the difference between the reconstructed and original features) to remove residual site effects, which are subsequently added back to produce the final harmonized output.
Distribution differences in covariates (e.g., age and sex) are common and often unavoidable in multi-site datasets. As theoretically demonstrated by Tachet et al. [111], directly applying cVAE under such conditions may lead to covariate-driven variations being incorrectly attributed to site effects. To address this issue, DeepResBat [5] introduces a two-stage strategy. Specifically, it first estimates covariate influences using nonlinear regression models. The covariate residuals, obtained by subtracting the estimated covariate contributions from the original features, are then used as input to a cVAE model to isolate and remove site-specific effects, yielding harmonized residuals. The final harmonized features can be reconstructed by reintroducing the covariate effects into the harmonized residuals (Figure 4). By targeting residuals rather than raw features for deep learning harmonization, DeepResBat explicitly preserves biological variability while effectively reduces the risk of spurious associations.
Figure 4: Schematic illustration of the feature-level deep learning harmonization method DeepResBat [5]. The covariates effect of the original features were first removed by subtraction and used as the input to a VAE. Then, the VAE output was added back to the removed covariate components to obtain harmonized features. ROI: region of interest.
5.3 Modality applicability and limitations
Feature-level methods operate on MRI-derived measurements or features rather than raw image intensities, and therefore are applicable to nearly all MRI modalities and are particularly well suited for large-scale, existing multi-center MRI datasets. Their flexibility, relatively strong generalizability, interpretability, and modest computational requirements have contributed to their continued status as the most widely adopted harmonization paradigm to date. As discussed above, ComBat and its extensions have been successfully applied to a broad range of feature types, including cortical thickness (structural MRI) [42], fractional anisotropy (dMRI) [43], functional connectivity (fMRI) [141], and even metabolite concentrations (MR spectroscopy) [14]. Owing to their algorithmic simplicity and practical feasibility, feature-level harmonization methods have been supported by well-established implementations, including ComBatHarmonization [^2] https://github.com/Jfortin1/ComBatHarmonization/, neuroHarmonize [^3] https://github.com/rpomponio/neuroHarmonize, and DPABI harmonization [^4] https://rfmri.org/content/dpabi-harmonization-toolbox-harmonizing-multi-site-brain-imaging-big-data-era, thereby facilitating their broad adoption in large-scale and translational studies.
Feature-level methods rely on a set of statistical assumptions, including linear or parametric formulations of site effects and the assumption that site-effect parameters across features are drawn from shared prior distributions [43, 26, 5]. These assumptions tend to underestimate spatially varying site effects, which in MRI are commonly introduced by B0 and B1 inhomogeneities as well as gradient nonlinearity. In addition, feature-level approaches are highly dependent on feature definitions and upstream processing pipelines. Systematic biases introduced during preprocessing cannot be corrected by subsequent harmonization, and as a direct consequence, derived features are often not comparable across studies employing different processing strategies. Compared with image-level approaches, feature-level methods do not produce harmonized images and are inherently tied to specific features and downstream tasks.
6 Traveling Subject
Harmonization models based on non-traveling subject datasets, whether traditional statistical approaches or learning-based methods, can effectively eliminate site effects. However, it remains unclear whether such models may also overcorrect biological variability. In this context, traveling subject-based approaches offer a baseline for rigorously disentangling biological and non-biological sources of variability. Moreover, publicly available traveling subject datasets not only enable investigation into how site effects influence multi-site statistical analyses, but also serve as valuable benchmarks for validating newly proposed harmonization methods.
6.1 Statistical approaches
Building on image-based histogram matching methods, Wrobel et al. proposed Multisite Image Harmonization by Cumulative Distribution Function Alignment (MICA) [134], which performs image harmonization based on the alignment of CDFs. The method first applies preprocessing steps such as N4 bias field correction and skull stripping to the images, and then computes their CDFs. For each traveling subject, MICA selects one image as the template and uses its CDF as the alignment target. It then estimates a nonlinear, monotonically increasing warping function by densely sampling paired points between the source and template images and applying linear interpolation. This warping function is used to align the CDF of the source image to that of the template.
Based on traveling subject (TS) data, ComBat can also be extended to the TS-ComBat method [87]. In this approach, the covariate term in the standard ComBat model is replaced with individual effects estimated from traveling subjects, while the site effects are still estimated and removed using an empirical Bayes method. To further account for repeated measurements across time points, Beer et al. [12] applied Longitudinal ComBat in the context of traveling subject studies. The model is expressed as follows:
$y_{ijv}(t)=a_{v}+\mathbf{X}_{ij}(t)\beta_{v}+\eta_{jv}+\gamma_{iv}+\delta_{iv}\varepsilon_{ijv}(t)$ (4)
where both $y_{ijv}(t)$ and $\varepsilon_{ijv}(t)$ introduced time-varying dependent variables, $\eta_{jv}$ represents subject-specific random intercept. In addition, several of the previously discussed image-level and feature-level harmonization methods have been further validated and extended on traveling subject datasets, demonstrating their adaptability across sites [87, 34, 138].
6.2 Learning-based approaches
Traveling-subject data naturally provide paired training samples for learning-based methods, offering a more direct solution compared to unpaired approaches. Methodologically, these techniques can be categorized into two main types: end-to-end mapping and anatomy-contrast disentanglement strategies similar to those used in unpaired settings (Figure 5).
Figure 5: Two representative deep learning-based harmonization strategies using traveling subject data: (a) end-to-end mapping and (b) anatomy-contrast disentanglement methods. The availability of paired training data provides additional supervision related to site or subject identity, which enhances the learning of site-invariant representations. A: anatomy; S: style.
6.2.1 End-to-end mapping
Based on paired training data, Dewey et al. proposed DeepHarmony [37], a U-Net-based harmonization framework. In their study, 12 subjects were scanned on two different scanners using protocols (e.g., T1, T2, PD, and FLAIR) with varying parameters. U-Net was then directly trained to learn an image-to-image mapping between paired images acquired from the two sites. DeepHarmony was shown to substantially reduce inter-protocol volumetric discrepancies in longitudinal MRI datasets of patients with multiple sclerosis.
Unlike methods that perform direct mapping in the image domain, Tong et al. [119] proposed a harmonization approach that maps source DW images to target diffusion kurtosis imaging (DKI) parameters. A 3D hierarchical convolutional neural network was trained using co-registered labels estimated through an iteratively reweighted linear least squares method. This approach resulted in a 50-60% reduction in inter-scanner variation of DKI parameters within white matter. Similarly, Tax et al. [114] and Ning et al. [95] summarized several learning-based harmonization methods from the Multi-Shell Diffusion MRI Harmonization Challenge (MUSHAC). These methods harmonize DWI data in the spherical harmonics domain. All included methods significantly reduced variability across multi-scanner DWI acquisitions, although challenges remain in accurately capturing localized features.
To improve generalizability to unseen sites, Xu et al. proposed Site Mix (SiMix) [136], which combines mixed-site training with test-time perturbation. Instead of harmonizing to a single existing site, SiMix constructs a virtual target site by linearly combining images from multiple known sites during training. At inference, the test image is mixed with its initially harmonized output to generate multiple perturbed inputs, whose predictions are averaged to produce the final harmonized result, following an ensemble strategy.
6.2.2 Anatomy-contrast disentanglement
Compared to non-traveling-subject methods, traveling-subject-based anatomy-contrast disentanglement offers the distinct advantage of efficiently utilizing shared anatomical structures across different sites for strong supervision, thereby enabling more accurate interpretation and quantification of site and scanner effects (Figure 5b).
A representative method is Multi-scanner Image harmonization via Structure Preserving Embedding Learning (MISPEL), proposed by Torbati et al. [122] The framework consists of scanner-specific encoders and decoders and follows a two-stage training strategy. This approach demonstrates that paired multi-site data can provide strong supervision, enabling the model to maintain high anatomical fidelity during harmonization. Notably, ESPA, proposed by Torbati et al. [123] and built on the MISPEL framework, relaxes the requirement for traveling-subject paired data by employing augmentation strategies on single-site images. Additionally, Tian et al. [116] proposed a bidirectional framework called deep learning-based representation disentanglement (DeRed). This framework consists of four encoders to disentangle anatomical and site-specific representations from paired different sites, and two decoders to bidirectionally synthesize harmonized images. A key advantage of this model is its flexible multi-site harmonization capability, where new unseen sites can be linked to the target site via intermediate domains without retraining the entire model.
6.3 Available traveling subject datasets
One of the major challenges in image harmonization is how to effectively evaluate its performance. A direct and reliable approach is to use traveling-subject datasets, which minimize the bias introduced by inter-site population sampling. However, acquiring such datasets is often costly and limited by the number of available participants. Therefore, leveraging existing publicly available traveling-subject datasets is often helpful. Table 2 summarizes the currently available public datasets, covering traveling subjects across different imaging modalities and age groups, and involving major scanner vendors or varying acquisition protocols [114, 130, 120, 113].
Table 2: Available traveling subject dataset
Dataset name Number of subjects (age) Number of Scanners/sites MRI modalities Data Repository
ON-Harmony [130] 20 (18-55y) 6/5 T1w, T2w, SWI, dMRI, rs-fMRI https://openneuro.org/datasets/ds004712
SRPBS [113] 9 (24–32y) 12/8 T1w, rs-fMRI, fieldmap https://bicr-resource.atr.jp/srpbsts
SDSU-TS [56] 9 (22-55 y) 2/2 T1w, T2w, dMRI https://openneuro.org/datasets/ds005664
HAMLET 5 (N/R) 4/3 T1w, dMRI, rs-fMRI https://www.nitrc.org/projects/hamlet
ZJU dMRI [120] 3 (23-26 y) 10/10 T1w, dMRI https://doi.org/10.6084/m9.figshare.8851955.v6
SPINS Human Phantoms [57] 4 (N/R) 6/3 T1w, dMRI, rs-fMRI https://openneuro.org/datasets/ds003011
MUSHAC [114] 14 (21-41 y) 3/ (N/R) dMRI https://www.cardiff.ac.uk/cardiff-university-brain-research-imaging-centre/research/projects/cross-scanner-and-cross-protocol-diffusion-MRI-data-harmonisation
N/R: Not reported
Taking ON-Harmony [130] as an example, 20 healthy volunteers were scanned using five imaging modalities across six scanners from different vendors and models. These modalities included structural imaging (T1-weighted, T2-weighted, and susceptibility-weighted imaging) as well as functional imaging (dMRI and resting-state fMRI). As shown in Figure 6, a clear observation is that functional modalities exhibit substantially greater inter-scanner variability than structural ones. This discrepancy arises not only from differences in reconstruction and post-processing pipelines across scanners, but also from the fact that both dMRI and fMRI typically rely on fast echo-planar imaging sequences for data acquisition, which are more susceptible to imperfections such as field inhomogeneities.
Figure 6: Representative examples of different modalities for a single participant data across all scanners from ON-Harmony dataset [130].
7 Evaluation Metrics
The challenge of validating harmonization remains a central bottleneck in the field. The absence of a definitive ground truth, together with the limitations of existing evaluation metrics, often renders validation even more challenging than harmonization itself. As a result, objective method comparison becomes challenging, and clinical translation is hindered. In this section, we review five categories of validation strategies reported in the literature. A comparative synthesis of MRI harmonization method families and their corresponding validation paradigms is summarized in Table 3. It should be noted that none of these strategies alone can conclusively establish biological fidelity. Instead, they assess different aspects of harmonization effectiveness through complementary yet inherently limited evaluation perspectives.
Table 3: Comparative synthesis of MRI harmonization method families
Harmonization level Primary object of harmonization Advantages Weaknesses Typical validation
Acquisition or reconstruction-level (Section 3) Signal formation and reconstruction • Controls variability at the signal source • High interpretability • Not applicable to retrospective datasets • Depends on vendor support • Challenging for large-scale deployment • Traveling subjects • Phantoms
Image-level (traditional) (Section 4.1) Image intensity or signal representation statistics • Simple and efficient • No training required • Broad applicability • Simple intensity transformations • Limited for complex site effects • Risk of removing biological variability • Traveling subjects • Visual assessment • Distributional statistics • Site discriminability tests • Downstream tasks performance
Image-level (learning-based) (Section 4.2) Image appearance or contrast • Strong visual harmonization • Models complex site effects • Limited interpretability • Weak biological constraints • Risk of over-correction or hallucination • Data intensive
Feature-level (statistical) (Section 5.1) Image-derived feature distributions • Explicit Biological modeling • Statistically interpretable • Well established • Relies on simple assumptions • Limited capacity for nonlinear or high-dimensional site effects • Sensitive to feature definition and sample imbalance • No harmonized images • Biological association analysis • Robustness tests under confounding • Site discriminability tests • Downstream tasks performance
Feature-level (learning-based) (Section 5.2) Learned feature representations • Flexible nonlinear modeling • Improved site-feature disentanglement • Reduced interpretability • Limited cross-task transferability
Traveling-subject -based (Section 6) Paired inter-site differences • Direct removal of site effects • Minimal modeling assumptions • Requires traveling subjects • Limited scalability • Traveling subjects
7.1 Reference-based evaluation
Reference-based evaluation is generally regarded as the most direct and interpretable strategy for assessing harmonization performance, as the underlying biological or physical state can be reasonably assumed to remain unchanged. Common forms include traveling subjects and phantoms. Such evaluation is not limited to the retrospective use of traveling-subject datasets described in Section 6, but also encompasses prospective validation in harmonized data acquisition frameworks (Section 3).
For image-based harmonization methods, traveling subjects enable direct voxel-wise comparison between harmonized images, allowing the use of quantitative image similarity metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Mean Absolute Error (MAE) [78, 22, 147]. In the case of feature-level harmonization, features extracted after harmonization can be evaluated using paired statistical tests (e.g., paired t-tests), Bland–Altman analysis, intra-class correlation coefficients (ICC), and coefficients of variation (CoV), thereby quantifying cross-site consistency and reliability. Compared with traveling subjects, phantoms provide a more controllable and repeatable imaging object, while their ability to represent true biological variability is inherently limited. Their standardized manufacturing and calibration protocols, such as those established by the National Institute of Standards and Technology (NIST), allow phantom-based validation to be performed across multiple sites or centers without requiring traveling scans. These phantoms have been widely used to assess the consistency of quantitative parameters in harmonized acquisition, particularly in qMRI and diffusion MRI [45, 82, 48, 70].
7.2 Reduction of site-related variability
Although existing studies employ a seemingly diverse set of statistical tests and classification experiments, their underlying objective is largely the same: to determine whether detectable site-related differences remain in imaging features after harmonization. In current practice, this objective is typically addressed through two complementary evaluation strategies.
Univariate testing is commonly used to assess whether individual imaging-derived measurements, such as ROI-level summary features or voxel-wise intensities, remain significantly associated with sites [43, 42, 44]. These approaches are based on the assumption that, if harmonization is successful, individual features should no longer exhibit systematic differences across sites. A typical implementation treats site as a categorical factor in ANOVA or linear models to quantify residual mean shifts between sites. Related tests, including Bartlett’s or Levene’s tests, focus instead on variance or scaling differences, and are used to evaluate whether harmonization also corrects site-specific differences in noise magnitude or dispersion. More general distributional differences beyond mean and variance can also be assessed using nonparametric tests such as the Kolmogorov-Smirnov test [33, 133].
A second widely adopted strategy is site discriminability testing, in which harmonized features are used as inputs to train classifiers (support vector machines or XGBoost) to predict the acquisition site [43, 5, 63, 6]. Within this framework, a reduction in site classification accuracy is interpreted as evidence of successful harmonization. Compared with univariate hypothesis testing, site discriminability provides a stronger assessment of residual site signatures in high-dimensional feature spaces. However, it is important to note that reduced site predictability does not necessarily imply adequate preservation of biological signals, as excessive harmonization may also remove meaningful biological variability while still achieving low site classification performance.
7.3 Preservation of biological variability
Compared with evaluating the reduction of site effects, verifying the preservation of biologically meaningful variability is more challenging. Although existing evaluations remain incomplete, prior studies have made several efforts in this direction, including association analyses with biological variables, robustness tests under confounding, biological variable prediction performance, and false-positive or permutation-based sanity checks.
Among existing evaluation strategies, biological variable association analysis is the most widely used, encompassing both univariate and multivariate approaches [43, 5, 63, 6]. The core question addressed by this strategy is whether statistical relationships between imaging features and known biological variables (e.g., age, sex, or clinical diagnosis) are preserved after harmonization. In practice, these relationships are typically assessed using generalized linear models (GLM) or ANOVA. Commonly reported metrics include regression coefficients or effect size together with their corresponding z-statistics or p-values. Beyond testing statistical significance, some studies further quantify biological preservation by examining the proportion of variance in imaging features explained by biological covariates. This is commonly implemented using linear regression models, where metrics such as $R^{2}$ or partial $R^{2}$ are used to measure the fraction of variability attributable to biological factors. A representative example is provided by ComBat [42], which evaluates the linear association between average cortical thickness and age before and after harmonization, using changes in $R^{2}$ as a quantitative indicator of age-related variance preservation.
Confounding between biological variables and site effects poses an additional challenge for MRI harmonization, for example when the distribution of age is imbalanced across sites. If not carefully addressed or explicitly modeled, harmonization procedures may inadvertently remove biologically meaningful age-related variability while attempting to eliminate site-related effects. One widely adopted strategy is to explicitly construct datasets with known confounding structures to evaluate harmonization robustness [43]. For example, both positively and negatively confounded scenarios can be derived from existing data, or predefined biologically associated and null voxels/ROIs can be used to compare effect sizes before and after harmonization. These analyses assess whether improved sensitivity to true biological effects is achieved without compromising specificity through spurious signal amplification.
Biological variable prediction performance provides another task-oriented and indirect assessment and is discussed in Section 7.4. Notably, for voxel-wise diffusion MRI features, prediction performance is often not the preferred evaluation metric due to the extremely high dimensionality of the data and the presence of well-established biological priors (e.g., known age-related trends in FA). Instead, prediction-based evaluation is often replaced by effect size consistency analysis, which examines whether biological effects estimated from harmonized pooled data are consistent with site-wise estimates in terms of spatial distribution, direction, and magnitude [43].
Finally, given that harmonization procedures may artificially amplify apparent biological associations even in the absence of true biological differences, some studies further incorporate false positive rate assessments or permutation-based sanity checks [5]. The central idea is to repeat association or prediction analyses after randomly permuting biological labels, if harmonized data continue to exhibit statistically significant associations or prediction performance exceeding chance levels, this suggests that the method may have introduced spurious biological signals. Recent study has demonstrated that such strategy constitutes an effective and practical approach for detecting hallucination effects in learning-based methods [5].
7.4 Performance on downstream tasks
Downstream tasks provide an additional important dimension for evaluating the harmonization methods in real-world applications. Such evaluations commonly rely on multi-fold cross-validation to fully leverage available data, while also assessing model generalizability under cross-site or cross-dataset validation. Existing studies can be broadly categorized into three types:
  1. 1.
    Classification: Representative tasks include the discrimination between individuals with autism spectrum disorder and healthy controls, as well as a range of Alzheimer’s disease-related classification problems. These tasks are commonly used to assess whether harmonization helps mitigate site-related effects in disease classification. Typical evaluation metrics include the area under the receiver operating characteristic curve (AUC), classification accuracy (ACC), sensitivity (SEN), and specificity (SPE) [22, 51] .
  2. 2.
    Segmentation: Typical tasks include brain tissue segmentation (e.g., gray matter, white matter, and cerebrospinal fluid) as well as white matter lesion segmentation in multiple sclerosis. Such tasks are most frequently conducted in structural MRI (e.g., T1- and T2-weighted imaging) and are primarily used in image-based harmonization. Common quantitative metrics include the Dice Similarity Coefficient (DSC) and the 95th percentile Hausdorff Distance (HD95) [13, 22, 135, 147].
  3. 3.
    Regression: A representative application is age prediction based on whole-brain signals or imaging-derived features (e.g., cortical thickness), which is used to assess the extent to which harmonization methods preserve continuous biological gradients. Common evaluation metrics include the root mean squared error (RMSE), mean squared error (MSE), and MAE [42, 147, 10].
However, it is important to note that improvements in downstream task performance do not necessarily imply biologically faithful or valid harmonization. Performance gains may arise from effective suppression of site-related variability, but may also result from over-correction or the attenuation of biologically meaningful information that is irrelevant to the specific task.
7.5 Qualitative and visual assessment
Qualitative visualization strategies for harmonization evaluation can generally be categorized into two classes: feature-level and image-level visualizations. Feature-level visualization focuses on examining the distributions of extracted features across different sites under matched or comparable biological conditions. Typical approaches include visualizing normalized intensity histograms[107, 134]. For harmonized features or learned latent representations, dimensionality reduction techniques such as Principal Component Analysis (PCA) and t-distributed stochastic neighbor embedding (t-SNE) are commonly used to visualize site-related separability before and after harmonization. In addition, some learning-based methods explicitly learn low-dimensional representations during model training, enabling direct visualization of how samples from different sites are geometrically aligned in the representation space following harmonization [147, 135, 146, 22].
In contrast, image-level visualization is primarily adopted in image-based harmonization and harmonized acquisition studies, where evaluation relies on direct visual inspection of contrast consistency and style differences across sites. Such visual assessment remains particularly important in pathology-sensitive scenarios, as it provides one of the few practical means to assess hallucination risk, and to identify spurious structures or anatomically implausible alterations introduced by harmonization procedures [13].
8 Discussion
8.1 Image-level or feature-level harmonization ?
In retrospective harmonization, the choice between image-level and feature-level methods is primarily determined by the harmonization objective and the characteristics of the available datasets.
Image-level harmonization is most appropriate when the goal is to produce usable harmonized images, for example, in clinical workflows, expert visual assessment, or downstream analyzes that are not yet defined at the time of harmonization. In these scenarios, task-agnostic correction enables broad reuse of the harmonized data, supports subsequent processing steps such as segmentation, classification, lesion visualization, and facilitates long-term data value and public dataset construction [24]. Moreover, image-level approaches offer the potential to mitigate spatially non-uniform or anatomically coupled site effects that cannot be adequately modeled at the feature level. When paired data are available, such as multi-contrast acquisitions or traveling-subject designs, image-level methods can exploit these constraints to better preserve anatomical consistency [122, 136, 146]. However, in the absence of paired data, image-level harmonization becomes substantially more challenging and typically relies on weaker implicit assumptions, increasing the risk of bias.
Feature-level harmonization remains the dominant choice in current retrospective studies, largely represented by ComBat and their extensions [43, 42]. These methods are particularly well suited for hypothesis-driven studies with clearly defined analysis targets, where downstream tasks and imaging features are fixed and statistical inference is the primary objective [145]. Feature-level approaches are also preferable when sample sizes are limited or site distributions are highly imbalanced, conditions under which image-level learning-based methods are prone to learning site-specific artifacts, introducing reference-site bias, or generating hallucinated effects that are difficult to detect. In addition, feature-level methods allow explicit modeling and control of biological covariates, which is critical when site effects are partially confounded with age, sex, or clinical variables. Within this category, traditional statistical models provide stable and interpretable correction in the absence of paired data, whereas learning-based feature-level methods, such as DeepResBat [5], offer increased flexibility in modeling nonlinear site effects while still retaining explicit covariate control.
Overall, while image-level harmonization represents an important long-term goal, feature-level methods currently offer a more reliable and interpretable solution for most retrospective multi-site studies. Image-level approaches should therefore be considered primarily in settings where harmonized images are explicitly required and sufficient data constraints are available to support rigorous validation.
8.2 Clinical translation and deployment challenges
While harmonization methods are currently most widely used in research settings for multi-center and large-scale studies, there are additional considerations for clinical deployment. Regulatory approval can be a major barrier in clinical deployment of image harmonization methods, especially those based on machine learning [112]. Regulators such as the FDA (Food and Drug Administration) and EMA (European Medicines Agency) require clear evidence that a harmonization method does not alter clinically relevant information or introduce systematic bias that could affect diagnosis, longitudinal follow-up, or treatment decisions. Demonstrating robust performance across platforms and software versions may require extensive validation. Both FDA [^5] https://www.fda.gov/regulatory-information/search-fda-guidance-documents/marketing-submission-recommendations-predetermined-change-control-plan-artificial-intelligence and EMA [^6] https://www.ema.europa.eu/en/news/reflection-paper-use-artificial-intelligence-lifecycle-medicines have provided new rules and guidelines for continuous monitoring of methods that may change over time, such as those based on continuous learning and federated learning.
Widespread clinical deployment of harmonization methods may also require vendor support. While developers or users (e.g., hospitals) can potentially maintain custom software, workflows, and ensure security and compliance on their own, image harmonization may be best achieved if it is integrated with vendor reconstruction pipelines. This may not be possible without vendor support or involvement.
8.3 Future Directions
With increasing efforts to harmonize the MRI workflow across the full acquisition-to-analysis pipeline, and driven in particular by advances in deep learning, the field is poised to benefit from several emerging opportunities :
  • 1.
    Harmonized acquisition and reconstruction: Previous research has primarily focused on retrospective harmonization, whereas harmonized acquisition and reconstruction strategies have received comparatively less attention [62, 1, 97]. As an approach that minimizes variability at the source, this strategy is undoubtedly one of the promising directions for future development. Further progress will depend on simplifying implementation and promoting the sharing of standardized acquisition and reconstruction workflows. When combined with retrospective harmonization approaches, these strategies have the potential to achieve more effective and robust harmonization across sites, while jointly optimizing the overall MRI acquisition pipeline in both clinical and research settings.
  • 2.
    Establishing standardized validation benchmarks: There has long been an urgent need to establish unified validation benchmarks that are quantitative and directly comparable across studies. The public release of large-scale, multimodal traveling-subject datasets, such as ON-Harmony[130], now provides a critical opportunity to advance this effort. Such benchmarks are particularly important for addressing the current landscape of harmonization research, especially the rapid proliferation of deep learning-based methods, for which evaluation protocols remain heterogeneous. At the same time, existing evaluation frameworks themselves remain limited and call for further methodological innovation, including the quantitative assessment of image-level hallucinations, as well as the evaluation of uncertainty and newly introduced biases associated with learning-based approaches.
  • 3.
    Integrating statistical and deep learning approaches: While deep learning offers powerful nonlinear modeling capabilities at both the image and feature levels, existing evidence suggests that such models may inadvertently suppress biologically meaningful signals, even when biological covariates are explicitly modeled[5]. DeepResBat [5] provides a representative example at the feature level, demonstrating that deep learning-based residual modeling can improve harmonization performance while still requiring careful control to avoid overcorrection of biological variability. In contrast, comparable image-level harmonization approaches that explicitly address this trade-off remain largely unexplored. From a methodological perspective, statistical approaches tend to be more robust in small-sample regimes or under severe confounding, whereas deep learning is capable of modeling complex spatial and structural patterns in images. Integrating these complementary strengths may therefore represent a promising direction for developing harmonization methods that are both flexible and biologically faithful.
  • 4.
    Data privacy and security: Traditional harmonization methods typically rely on centralized data processing, which is often impractical given the sensitive nature of medical images and cross-institutional data-sharing constraints. Federated learning offers an alternative by enabling local model training with only parameter sharing, thereby, in principle, alleviating the need for centralized harmonization [52, 80]. However, adapting existing harmonization strategies to federated settings remains challenging, as many methods assume joint access to source and target domains or paired training data, conditions that are rarely met in distributed multi-site studies [80]. Beyond federated learning, an emerging complementary direction is to explicitly model and simulate unwanted sources of variability during training, as exemplified by synthetic data-driven approaches [140, 139, 129, 66, 16]. By exposing models to controlled, diverse forms of site-related bias, such strategies aim to improve robustness and reduce data-sharing constraints, offering a promising alternative for privacy-preserving harmonization.
In summary, harmonization as a preprocessing step for large-scale MR image analysis continues to offer substantial opportunities and plays an important role that is difficult to replace. At the same time, as these methods are increasingly applied and further developed, their limitations and potential risks should be carefully acknowledged. Method selection should therefore be guided by specific analytical goals and supported by rigorous and comprehensive validation, to ensure appropriate and responsible use of harmonization in practice.
9 Conclusion
This survey reviews recent advances in MRI harmonization across the full imaging pipeline, spanning harmonized image acquisition, retrospective image-level and feature-level methods, evaluation strategies, and publicly available traveling-subject datasets. Nevertheless, although substantial progress has been made in reducing site effects, solid evidence for biological signal preservation and standardized validation benchmarks are still lacking. Accordingly, this survey provides a comprehensive reference for the field and highlights key methodological gaps and future directions toward developing reliable, biologically informed harmonization frameworks for multi-site MRI studies.
CRediT authorship contribution statement
Qinqin Yang: Writing - review & editing, Writing - original draft, Visualization, Resources, Methodology, Investigation, Conceptualization. Firoozeh Shomal-Zadeh: Writing - review & editing, Writing - original draft, Visualization, Investigation. Ali Gholipour: Supervision, Conceptualization, Writing - review & editing, Validation, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This research was supported in part by the National Institute of Health (NIH) under award numbers R01EB031849, R01EB032366, and R01HD109395. The content of this publication is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Data availability
No data was used for the research described in the article.